### v1

Abstract
We analyze the Helmholtz–Korteweg and nematic Helmholtz–Korteweg equations, variants of the classical Helmholtz equation for time-harmonic wave propagation for Korteweg and nematic Korteweg fluids. Korteweg fluids are ones where the stress tensor depends on density gradients; nematic Korteweg fluids further depend on a nematic director describing the orientation of the (anisotropic) molecules. We prove existence and uniqueness of solutions to these equations for suitable (nonresonant) wave numbers and propose convergent discretizations for their numerical solution. The discretization of these equations is nontrivial as they demand high regularity and involve unfamiliar boundary conditions; we address these challenges by using high-order conforming finite elements, and enforcing the boundary conditions with Nitsche’s method. We illustrate our analysis with numerical simulations in two dimensions.
keywords:
Korteweg fluids, nematic liquid crystals, Helmholtz equation, wave propagation
{MSCcodes}
65M60, 35J05
1 Introduction
The propagation of time-harmonic acoustic waves in inviscid elastic fluids with stress tensor $\sigma=-pI$ ($p$ the pressure) is given by the classical Helmholtz equation [25]
(1) $-\Delta u-k^{2}u=f,$
where $u$ is the density of the fluid, $f$ is a time-harmonic forcing term, and $k$ is the wave number. If the stress tensor has additional terms, then the corresponding Helmholtz equation also has additional terms. In particular, if the stress tensor has a Korteweg term depending on density gradients [31, 38, 39]
(2) $\sigma=-pI-\alpha u\left(\nabla u\otimes\nabla u\right),$
where $\alpha>0$ is a constitutive parameter, then the resulting Helmholtz–Korteweg equation is given by
(3) $\alpha\Delta^{2}u-\Delta u-k^{2}u=f.$
Nematic liquid crystals are Korteweg fluids with an additional anisotropic term in the stress tensor that depends on the director $\mathbf{n}\in\mathbb{S}^{2}$ [42]:
(4) $\sigma=-pI-\alpha u\left(\nabla u\otimes\nabla u\right)-\beta\left(\nabla u\cdot\mathbf{n}\right)\nabla u\otimes\mathbf{n},$
where $\beta>0$ is a constitutive parameter. Under the assumption that $\nabla\mathbf{n}$ vanishes at the acoustic length-scale, the resulting nematic Helmholtz–Korteweg equation is given by
(5) $\alpha\Delta^{2}u+\beta\nabla\cdot\nabla\left(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\right)-\Delta u-k^{2}u=f,$
where $\mathcal{H}u$ is the Hessian matrix of $u$. Here $\mathbf{n}$ describes the local average orientation of the (calamitic, i.e. rodlike) molecules in the fluid. The Helmholtz–Korteweg and nematic Helmholtz–Korteweg equations were recently derived by a subset of the authors [18]. The solutions of the nematic Helmholtz–Korteweg equation exhibit surprising phenomena not permitted by (1), including anisotropy in propagation of sound, penetration depth, and scattering [18].
In this manuscript we consider the PDE analysis and numerical analysis of the Helmholtz–Korteweg and nematic Helmholtz–Korteweg equations (3) and (5). We prove existence and uniqueness of solutions for suitable (nonresonant) $k$. While the numerical analysis of (1) has been intensively studied [26, 17, 32, 11], (3) and (5) present new challenges. In particular, the biharmonic and Hessian terms demand the discretization of $u\in H^{2}$, and the various boundary conditions are difficult to enforce strongly or naturally. We overcome these challenges by using high-order $H^{2}$-conforming finite elements, and by enforcing the boundary conditions with Nitsche’s method. Conforming $H^{2}$ discretizations are naturally high-order, which is sometimes seen as a disadvantage; in this context it is appealing, since high-order discretizations achieve quasi-optimality for large wave numbers on much coarser meshes [19, 15, 28, 29].
We consider the boundary value problem
[…]
where $\Omega\subset\mathbb{R}^{d}$, $d\in\{2,3\}$, is either a bounded convex polytope or a bounded domain with smooth boundary, and $\mathcal{B}$ is an operator encoding the boundary conditions. In particular, $\mathcal{B}$ can encode sound soft, sound hard, or impedance boundary conditions. Sound soft boundary conditions impose that the acoustic pressure vanishes on the boundary, which corresponds to
(7) $\mathcal{B}u=(u,\Delta u+\frac{\beta}{\alpha}\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}).$
Sound hard boundary conditions impose that the normal derivative of the fluid velocity vanishes on the boundary, which corresponds to
(8) $\mathcal{B}u=(\partial_{n}u,\partial_{n}\Delta u+\frac{\beta}{\alpha}\partial_{n}(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n})).$
Lastly, impedance boundary conditions impose that the normal derivative of the fluid velocity is proportional to the fluid velocity, which corresponds to
(9) $\mathcal{B}u=(\partial_{n}u-i\theta u,\partial_{n}\Delta u-i\theta(\Delta u+\frac{\beta}{\alpha}\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}-\frac{\beta}{\alpha}\partial_{n}(\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n}))),$
for specified impedance $\theta$. We next prove the well-posedness of the continuous boundary value problem (6) for each type of boundary conditions introduced.
2 Weak formulation
Integration by parts of the first (biharmonic) and third (Laplacian) terms is standard. Integrating the second (nematic) term by parts twice yields
$\beta$ $\int_{\Omega}\nabla\cdot\nabla(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n})v\,\mathrm{d}x$
(10) $=\beta\!\int_{\Omega}\!\!\left(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\right)\Delta v\,\mathrm{d}x-\beta\!\int_{\partial\Omega}\!\!\left(\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n}\right)\nabla v\cdot\bm{\nu}\,\mathrm{d}s+\beta\!\int_{\partial\Omega}\nabla\left(\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n}\right)\cdot\bm{\nu}v\,\mathrm{d}s,$
where $\bm{\nu}$ is the outward-facing unit normal on $\partial\Omega$. The weak form of (6) is given by
(11) $\text{find }u\in X\text{ such that }a(u,v)=(f,v)_{L^{2}(\Omega)}\quad\forall v\in X,$
where we specify the bilinear form $a(\cdot,\cdot)$ and the space $X$ for each type of boundary condition.
For sound soft boundary conditions, we set $X=H^{2}_{0}(\Omega)\coloneqq H^{2}(\Omega)\cap H^{1}_{0}(\Omega)$ and define the bilinear form $a(\cdot,\cdot)$ to be
$a(u,v)\coloneqq$ $\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}\!-\!k^{2}(u,v)_{L^{2}(\Omega)},$
where the boundary terms vanish due to the boundary conditions. For sound hard boundary conditions, we set $X\coloneqq H^{2}(\Omega)$ and define the bilinear form $a(\cdot,\cdot)$ to be
(12) $a(u,v)\coloneqq$ $\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}-k^{2}(u,v)_{L^{2}(\Omega)}$
-- $\alpha(\Delta u,\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$
where the absent boundary terms vanish due to the boundary conditions. For impedance boundary conditions, we set $X\coloneqq H^{2}(\Omega)$ and define the sesquilinear form $a(\cdot,\cdot)$ to be
(13) $a(u,v)\coloneqq$ $\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}-k^{2}(u,v)_{L^{2}(\Omega)}$
-- $\alpha(\Delta u,\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)}+\alpha i\theta(\Delta u,v)_{L^{2}(\partial\Omega)}+\beta i\theta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},v)_{L^{2}(\partial\Omega)}$
-- $\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\nabla v\cdot\bm{n})_{L^{2}(\partial\Omega)}-i\theta(u,v)_{L^{2}(\partial\Omega)},$
where several boundary terms have been substituted using the boundary conditions.
3 Analysis of the Helmholtz–Korteweg equation ($\beta=0$)
To analyze the variational problem (11), we use the concept of (weak) T-coercivity [9, 14], which has originally been introduced to treat sign shifting coefficients [14]. It can also be used as an alternative to the inf-sup condition in a general setting, for example for the analysis of the Helmholtz-like problems [12, 41, 10] or Galbrun’s equation [22, 23, 21, 40] and their respective finite element discretizations.
Definition 3.1.
We call a sesquilinear form $a:X\times X\to\mathbb{C}$ T-coercive if there exists a bijective operator $T\in\mathcal{L}(X)$ and a constant $\gamma>0$ such that
(14) $\Re\{a(Tu,u)\}\geq\gamma\|u\|^{2}_{X}\quad\forall u\in X.$
Equivalently, we require that $AT$ is coercive, where $A\in\mathcal{L}(X)$ is the operator associated to the sesquilinear form $a(\cdot,\cdot)$.
It can be shown that T-coercivity is equivalent to the inf-sup condition [12] and thus, it is a necessary and sufficient condition for the well-posedness of the problem. Let us also introduce the notion of weak T-coercivity.
Definition 3.2 (weak T-coercivity).
An operator $A\in\mathcal{L}(X)$ is called weakly T-coercive if there exists a bijective operator $T\in\mathcal{L}(X)$ and a compact operator $K\in\mathcal{L}(X)$ such that $AT+K$ is bijective.
It is well-known that weakly T-coercive operators are Fredholm with index zero and therefore bijective if and only if they are injective [20, 40]. In the following, let us denote by $\mathcal{A}$ the operator induced by the bilinear (sesquilinear) forms defined in (2), (12), or (13), respectively. We split the operator into a T-coercive part $A$ and a compact part $K\coloneqq\mathcal{A}-A$, where $K$ varies depending on the boundary conditions. In particular, we define $A\in\mathcal{L}(X)$ as
(15) $(Au,v)_{H^{2}(\Omega)}\coloneqq\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}-k(u,v)_{L^{2}(\Omega)}.$
In the case of sound soft boundary conditions, we have that $K=0$.
To construct a $T$ operator that satisfies the desired properties, we consider the following self-adjoint eigenvalue problem: find $u\in H^{2}_{0}(\Omega)$, $\lambda\in\mathbb{C}$ such that
(16) $\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}=\lambda(u,v)_{L^{2}(\Omega)}\quad\forall v\in H^{2}_{0}(\Omega).$
Lemma 3.3.
The problem (16) is well-posed and the associated solution operator $S:L^{2}(\Omega)\to L^{2}(\Omega)$ is compact.
Proof 3.4.
This follows from Lemma 4.3 below with $\beta=0$.
A direct consequence of Lemma 3.3 is that the eigenvalues are real, of finite multiplicity and that the eigenfunctions form an orthonormal basis of $L^{2}(\Omega)$ [6]. We denote the eigenvalues by $\{\lambda^{(i)}\}_{i\in\mathbb{N}}$ and the corresponding eigenfunctions by $\{e^{(i)}\}_{i\in\mathbb{N}}$. Without loss of generality, we assume that the eigenfunctions are normalized with respect to the $H^{2}$-norm. From this point on we will assume that $k^{2}\not\in\{\lambda^{(i)}\}_{i\in\mathbb{N}}$. Let $i_{\ast}$ be the largest index such that $\lambda^{(i)}<k^{2}$, i.e. $i_{\ast}\coloneqq\max\{i\in\mathbb{N}:\lambda^{(i)}<k^{2}\}$. Then, we define [^1] For readability we neglect the trivial case where $k^{2}<\lambda^{(1)}$.
(17) $W\coloneqq\text{span}_{0\leq i\leq i_{\ast}}\{e^{(i)}\},\qquad T\coloneqq\operatorname{id}_{X}-2P_{W}$
where $P_{W}:X\to W$ is the orthogonal projection onto $W$, orthogonal in the bilinear form on the left-hand side of (16). Since $T^{2}=\operatorname{id}$, the operator $T$ is bijective.
Theorem 3.5.
The operator $A\in\mathcal{L}(X)$ defined by (15) is T-coercive with respect to the $T$-operator defined in (17).
Proof 3.6.
Let $u\in X$ be arbitrary and $u^{(i)}\in\mathbb{C}$ s.t. $u=\sum_{i\in\mathbb{N}}u^{(i)}e^{(i)}$. Then, we have
$(ATu,u)_{H^{2}}$ $=\sum_{0\leq i\leq i_{\ast}}C(\lambda^{(i)})(k^{2}-\lambda^{(i)})(u^{(i)})^{2}+\sum_{i>i_{\ast}}C(\lambda^{(i)})(\lambda^{(i)}-k^{2})(u^{(i)})^{2}$
$\geq\gamma\|u\|^{2}_{X}$
with $\gamma\coloneqq\min_{i\in\mathbb{N}}\{C(\lambda^{(i)})|\lambda^{(i)}-k^{2}|\}>0$, where $C(\lambda^{(i)})$ is the renormalization constant associated with the unitary condition of the eigenfunctions with respect to the $H^{2}$-norm. Thus, $A$ is indeed T-coercive.
In the case of sound soft boundary conditions, i.e. (2), this concludes the analysis of the continuous problem (11) as $K=0$. In the other cases, i.e. (12) and (13) we still have to show that the operator $K$ is compact and that the operator $\mathcal{A}$ is injective.
Lemma 3.7.
The operator $K\in\mathcal{L}(X)$ defined as $K=\mathcal{A}-A$, where $\mathcal{A}$ is the operator associated with (12) or (13), is compact.
Proof 3.8.
Let us consider first the case of sound hard boundary conditions. Let $\gamma_{0}:X\to L^{2}(\partial\Omega)$ denote the trace operator. We have that
$\|Ku\|_{H^{2}(\Omega)}=\sup_{v\in X\setminus\{0\}}\frac{|(Ku,v)_{H^{2}(\Omega)}|}{\|v\|_{H^{2}(\Omega)}}$ $\leq\sup_{v\in X\setminus\{0\}}\frac{|\alpha|\|\gamma_{0}\Delta u\|_{L^{2}(\partial\Omega)}\|\gamma_{0}\nabla v\cdot\bm{\nu}\|_{L^{2}(\partial\Omega)}}{\|v\|_{H^{2}(\Omega)}}$
$\leq C|\alpha|\|\gamma_{0}\Delta u\|_{L^{2}(\partial\Omega)},$
where we used the continuity of the normal trace operator.
Similarly, for impedance boundary conditions, we calculate
$\|Ku\|_{H^{2}(\Omega)}$ $=\sup_{v\in X\setminus\{0\}}\frac{|(Ku,v)_{H^{2}(\Omega)}|}{\|v\|_{H^{2}(\Omega)}}$
$\leq\sup_{v\in X\setminus\{0\}}\frac{|\alpha|\|\gamma_{0}\Delta u\|_{L^{2}(\partial\Omega)}(\|\gamma_{0}\nabla v\cdot\bm{\nu}\|_{L^{2}(\partial\Omega)}+|\theta|\|\gamma_{0}v\|_{L^{2}(\partial\Omega)})}{\lVert v\lVert_{H^{2}(\Omega)}}$
$+\sup_{v\in X\setminus\{0\}}\frac{|\theta|\|\gamma_{0}u\|_{L^{2}(\partial\Omega)}\|\gamma_{0}v\|_{L^{2}(\partial\Omega)}}{\|v\|_{H^{2}(\Omega)}}$
$\leq C(\|\gamma_{0}\Delta u\|_{L^{2}(\partial\Omega)}+\|\gamma_{0}u\|_{L^{2}(\partial\Omega)}),$
where we use the continuity of the normal trace operator and the compactness of the trace operator. Thus, for any sequence $(u_{n})_{n\in\mathbb{N}}$ such that $u_{n}\rightharpoonup u$ in $H^{2}$, it follows that $Ku_{n}\to Ku$ and therefore $K$ is compact.
Lemma 3.9.
The operator $\mathcal{A}$ associated with (12) or (13) is injective.
Proof 3.10.
For sound hard boundary conditions this follows from the assumption that $k^{2}\not\in\{\lambda^{(i)}\}_{i\in\mathbb{N}}$. For the impedance case, we essentially adapt the argument from the Helmholtz case, cf. [16, Thm. 35.5]. Taking $v\in\ker\mathcal{A}$ we obtain using Young’s inequality with $\zeta>0$
$0=|-\Im a(v,v)|=|-\alpha\theta(\Delta v,v)_{L^{2}(\partial\Omega)}+\theta\|v\|_{L^{2}(\partial\Omega)}|\geq\left|\frac{\alpha\zeta}{2}\|\Delta v\|^{2}_{L^{2}(\partial\Omega)}+\frac{\theta}{2\zeta}\|v\|_{L^{2}(\partial\Omega)}^{2}\right|$
which implies that $\gamma_{0}v=0$ and $\gamma_{0}\Delta v=0$ on $\partial\Omega$. Let $B_{R}$ be a ball of radius $R$ such that $\Omega\subset B_{R}$ which we assume to be centered at zero and let $v^{\mathcal{E}}_{R}$ be the extension by zero of $v$ onto $B_{R}\cap\Omega^{c}$. Then, we have that $v^{\mathcal{E}}_{R}\in H^{2}_{0}(B_{R})$ and since $\gamma_{0}\Delta v=0$ on $\partial\Omega$, we have that $\Delta v^{\mathcal{E}}_{R}=0$ in $B_{R}\cap\Omega^{c}$. Due to the Robin boundary condition, we have that $\partial_{n}\Delta v^{\mathcal{E}}_{R}|_{\partial\Omega}=0$. Hence $\nabla\Delta v$ is continuous across $\partial\Omega$. Since $(\nabla\Delta v^{\mathcal{E}}_{R})|_{\Omega}\in H(\operatorname{div},\Omega)$ and $(\nabla\Delta v^{\mathcal{E}}_{R})|_{B_{R}\cap\Omega^{c}}\in H(\operatorname{div},B_{R}\cap\Omega^{c})$, it follows that $\Delta^{2}v^{\mathcal{E}}_{R}\in L^{2}(B_{R})$. Since $v^{\mathcal{E}}_{R}$ is zero in an open subset of $B_{R}$, we can use the unique continuation principle for elliptic differential operators, cf. [27, Thm. 17.2.6], to conclude that $v^{\mathcal{E}}_{R}=0$ in $B_{R}$ and thus $v=0$ in $\Omega$.
4 Analysis of the nematic Helmholtz–Korteweg equation ($\beta>0$)
To extend the analysis from the previous section to the case where $\beta>0$, we consider the following modified eigenvalue problem: find $u\in H^{2}_{0}(\Omega)$, $\lambda\in\mathbb{C}$ such that
(18) $e(u,v)\coloneqq\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}=\lambda(u,v)_{L^{2}(\Omega)}$
for all $v\in H^{2}_{0}(\Omega)$. To employ the same argument as in the previous section, we have to show that the operator associated with (18) is self-adjoint.
Theorem 4.1.
The eigenvalue problem (18) is self-adjoint.
Proof 4.2.
Proving the self-adjointness of the operator associated with (18) is a matter of tedious integration by parts. We will omit most steps and only include salient ones. Let $u,v\in H^{2}_{0}(\Omega)$ be arbitrary. We begin observing that $\nabla\cdot\nabla[\mathbf{n}^{\top}(\mathbf{n}^{T}\nabla u)]$ can be expanded as
(19) $n_{x}^{2}\partial_{x}^{4}u+n_{y}^{2}\partial_{y}^{4}u+(n_{x}+n_{y})^{2}\partial_{x}^{2}\partial_{y}^{2}u+2(n_{x}n_{y})\partial_{x}^{3}\partial_{y}u+2(n_{x}n_{y})\partial_{x}\partial_{y}^{3}u,$
where we assumed that $\mathbf{n}=(n_{x},n_{y})^{\top}$, since by assumption we know that $\nabla\mathbf{n}$ is identically zero. By integration by parts, we immediately observe that the first three terms in (19) are self-adjoint. Due to the regularity of the functions $u,v\in H^{2}_{0}(\Omega)$, we can swap the order of differentiation and integrate by parts to show that the last two terms in (19) are self-adjoint as well.
Lemma 4.3.
For $\beta$ sufficiently small, the eigenvalue problem (18) is well-posed and the associated solution operator $S:L^{2}(\Omega)\to L^{2}(\Omega)$ is compact.
Proof 4.4.
Since $\|\mathbf{n}\|_{L^{\infty}(\Omega)}=1$ and $\beta>0$, we have that
$|\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta u)_{L^{2}(\Omega)}|\leq\beta\|\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n}\|_{L^{2}(\Omega)}\|\Delta u\|_{L^{2}(\Omega)}\leq\beta\|u\|^{2}_{H^{2}(\Omega)}.$
The Poincaré inequality yields for $\beta<\min\{\alpha,1/2,C_{P}^{2}/2\}$ that
$|e(u,u)|$ $=\alpha\|\Delta u\|^{2}_{L^{2}(\Omega)}+|\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta u)_{L^{2}(\Omega)}|+\|\nabla u\|^{2}_{L^{2}(\Omega)}$
$\geq\alpha\|\Delta u\|^{2}_{L^{2}(\Omega)}-\beta\|u\|^{2}_{H^{2}(\Omega)}+\frac{1}{2}\|\nabla u\|^{2}_{L^{2}(\Omega)}+\frac{C_{P}^{2}}{2}\|u\|^{2}_{L^{2}(\Omega)}$
$\geq\left(\min\{\alpha,1/2,C_{P}^{2}/2\}-\beta\right)\|u\|^{2}_{H^{2}(\Omega)}\geq\gamma_{*}\|u\|^{2}_{H^{2}(\Omega)},$
and hence $e(\cdot,\cdot)$ is coercive, with $\gamma_{*}\coloneqq\min\{\alpha,1/2,C_{P}^{2}/2\}-\beta>0$. The compactness of the solution operator follows from the compact embedding $H^{2}_{0}(\Omega)\hookrightarrow L^{2}(\Omega)$.
Due to the above results, we conclude that for sufficiently small $\beta$ the eigenvalues $\{\lambda^{(i)}\}_{i\in\mathbb{N}}$ of (18) are real, of finite multiplicity, and that the associated eigenfunctions $\{e^{(i)}\}_{i\in\mathbb{N}}$ form an orthogonal basis of $L^{2}(\Omega)$. As before, we assume that the eigenfunctions are normalized with respect to the $H^{2}$-norm.
We set $i_{\ast}\coloneqq\max\{i\in\mathbb{N}:\lambda^{(i)}<k^{2}\}$ and define
(20) $W\coloneqq\text{span}_{0\leq i\leq i_{\ast}}\{e^{(i)}\},\qquad T\coloneqq\operatorname{id}_{X}-2P_{W}.$
Furthermore, we redefine $A\in\mathcal{L}(X)$ as $(Au,v)_{H^{2}(\Omega)}\coloneqq e(u,v)-k^{2}(u,v)_{L^{2}(\Omega)}$ and $K\coloneqq\mathcal{A}-A$, where we recall that $K$ is different depending on the boundary condition.
Theorem 4.5.
The operator $A\in\mathcal{L}(X)$ is T-coercive with respect to the $T$-operator defined in (20).
Proof 4.6.
This follows from the same argument as in the proof of Theorem 3.5.
This concludes the analysis for the case of sound soft boundary conditions. For the remaining cases, we again show the compactness of the operator $K$ and the injectivity of the operator $\mathcal{A}$.
Lemma 4.7.
The operator $K\coloneqq\mathcal{A}-K\in\mathcal{L}(X)$ is compact.
Proof 4.8.
We estimate the new boundary terms arising from the nematic term. For sound hard boundary conditions, with the continuity of the normal trace operator we have
$|\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)}|$ $\leq\beta\|\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n}\|_{L^{2}(\partial\Omega)}\|\nabla v\cdot\bm{\nu}\|_{L^{2}(\partial\Omega)}$
$\leq C\beta\|\gamma_{0}\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\|_{L^{2}(\partial\Omega)}\|\nabla v\|_{L^{2}(\Omega)}.$
In the impedance case, we estimate
||| $\beta i\theta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},v)_{L^{2}(\partial\Omega)}-\beta(\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n},\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)}|$
$\leq C\beta(1+|\theta|)\|\gamma_{0}\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\|^{2}_{L^{2}(\partial\Omega)}\left(\|v\|_{L^{2}(\partial\Omega)}+\|\nabla v\|_{L^{2}(\Omega)}\right).$
Combining these estimates with the estimates from Lemma 3.7, we obtain for both cases that
$\|Ku\|_{H^{2}}\leq C\left(\|\gamma_{0}\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\|_{L^{2}(\partial\Omega)}+\|\gamma_{0}\Delta u\|_{L^{2}(\partial\Omega)}+\|\gamma_{0}u\|_{L^{2}(\partial\Omega)}\right).$
Since the trace operator $\gamma_{0}$ is compact, we have that $Ku_{n}\to Ku$ for any sequence $u_{n}\rightharpoonup u$ in $H^{2}$ and thus $K$ is compact.
Lemma 4.9.
The operator $\mathcal{A}\in\mathcal{L}(X)$ associated with (12) or (13) is injective.
Proof 4.10.
We extend the proof from Lemma 3.9 to account for the nematic term. Let $v\in\ker\mathcal{A}$. We have for $\zeta_{1},\zeta_{2}>0$ that
0= 0=|-0 = | - $\Im a(v,v)|=|-\alpha\theta(\Delta v,v)_{L^{2}(\partial\Omega)}-\beta\theta(\mathbf{n}^{\top}(\mathcal{H}v)\mathbf{n},\Delta v)_{L^{2}(\partial\Omega)}+\theta\|v\|_{L^{2}(\partial\Omega)}|$
$\geq\left|\frac{\alpha\zeta_{1}+\beta\theta\zeta_{2}}{2}\|\Delta v\|^{2}_{L^{2}(\partial\Omega)}+\frac{\beta\theta}{2\zeta_{2}}\|\mathbf{n}^{\top}(\mathcal{H}v)\mathbf{n}\|^{2}_{L^{2}(\partial\Omega)}+\frac{\theta}{2\zeta_{1}}\|v\|_{L^{2}(\partial\Omega)}^{2}\right|$
$\geq\left|\frac{\alpha\zeta_{1}+\beta\theta\zeta_{2}}{2}\|\Delta v\|^{2}_{L^{2}(\partial\Omega)}+\frac{\theta}{2\zeta}\|v\|_{L^{2}(\partial\Omega)}^{2}\right|.$
The rest of the proof follows analogously to the proof of Lemma 3.9.
5 Discretization of the Helmholtz–Korteweg equation ($\beta=0$)
In this section, we analyze the discretization of (11) using $H^{2}$-conforming finite elements. For brevity we focus our analysis on the sound soft and the impedance boundary conditions. We remark that sound hard boundary conditions can be analyzed similarly. Let $\{\mathcal{T}_{h}\}_{h}$ be a sequence of shape regular simplicial triangulations of $\Omega$. For polynomial degree $p>4$, we define the finite element space $X_{h}\subset H^{2}(\Omega)$ as
$X_{h}\coloneqq\{v\in H^{2}(\Omega):v|_{T}\in\mathcal{P}^{p}(T)\quad\forall T\in\mathcal{T}_{h}\},$
where $\mathcal{P}^{p}(T)$ is the space of polynomials of degree $p$ on an element $T\in\mathcal{T}_{h}$. To implement the boundary conditions in the sound soft case, we use Nitsche’s method [35]. To this end, we define the following terms:
(21a) $N^{\Delta^{2}}_{h}(u_{h},v_{h})$ $\coloneqq\alpha(\nabla(\Delta u_{h})\cdot\bm{\nu},v_{h})_{L^{2}(\partial\Omega)}+\alpha(u_{h},\nabla(\Delta v_{h})\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$
(21b) $N^{\Delta}_{h}(u_{h},v_{h})$ $\coloneqq(\nabla u_{h}\cdot\bm{\nu},v_{h})_{L^{2}(\partial\Omega)}+(u_{h},\nabla v_{h}\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$
(21c) $S_{h}(u_{h},v_{h})$ $\coloneqq\alpha\frac{\eta_{1}}{h^{3}}(u_{h},v_{h})_{L^{2}(\partial\Omega)}+\frac{\eta_{2}}{h}(u_{h},v_{h})_{L^{2}(\partial\Omega)},$
where $\eta_{1},\eta_{2}>0$ are suitably chosen stabilization parameters.
To define a discrete analogue of $T$, we consider the discretization of the eigenvalue problem (16): find $u_{h}\in\tilde{X}_{h}\subseteq X_{h}$, $\lambda\in\mathbb{R}$ s.t.
(22) $e_{h}(u_{h},v_{h})\coloneqq$ $\alpha(\Delta u_{h},\Delta v_{h})_{L^{2}(\Omega)}+(\nabla u_{h},\nabla v_{h})_{L^{2}(\Omega)}$
$+$ $\epsilon\left(N^{\Delta^{2}}_{h}(u_{h},v_{h})-N^{\Delta}_{h}(u_{h},v_{h})+S_{h}(u_{h},v_{h})\right)=\lambda(u_{h},v_{h})_{L^{2}(\Omega)}$
for all $v_{h}\in\tilde{X}_{h}\subseteq X_{h}$. In the forthcoming analysis, we set $\epsilon=1$ for the sound soft case and $\epsilon=0$ for the impedance case, where we assume for the latter case that the boundary conditions $u_{h}=0$ and $\Delta u_{h}=0$ on $\partial\Omega$ are imposed through the discrete space $\tilde{X}_{h}$ and naturally, respectively. For the sound soft case, we have that $\tilde{X}_{h}=X_{h}$ and the boundary conditions are imposed through the Nitsche terms (21).
Remark 5.1 (On Nitsche’s method).
The main motivation for using Nitsche’s method is the difficulty of imposing essential boundary conditions for $C^{1}$-conforming finite elements, cf. [30]. We note that the forthcoming analysis includes the case where this is achieved by setting $\epsilon=0$ and modifying $\tilde{X}_{h}$ also in the sound soft case.
5.1 Analysis of the discrete eigenvalue problem
Before considering the approximation of (11), we analyze the discrete eigenvalue problem (22). As a preparation, we show the following estimates.
Lemma 5.2.
There exist constants $C^{\Delta^{2}},C^{\Delta}>0$ s.t. for all $u_{h}\in X_{h}$ and all $\zeta_{1},\zeta_{2}>0$, we have the following estimates
$\alpha^{-1}N_{h}^{\Delta^{2}}(u_{h},u_{h})$ $\leq C^{\Delta^{2}}\zeta_{1}h^{-3}\|\Delta u_{h}\|^{2}_{L^{2}(\Omega)}+\zeta_{1}^{-1}\|u_{h}\|^{2}_{L^{2}(\partial\Omega)},$
$N_{h}^{\Delta}(u_{h},u_{h})$ $\leq C^{\Delta}\zeta_{2}h^{-1}\|\nabla u_{h}\|^{2}_{L^{2}(\Omega)}+\zeta_{2}^{-1}\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}.$
Proof 5.3.
First of all, we note that the following inequality holds [3, Lem. 5.2]:
$\|\nabla(\Delta u_{h})\cdot\bm{\nu}\|^{2}_{L^{2}(\partial\Omega)}\leq Ch^{-3}\|\Delta u_{h}\|^{2}_{L^{2}(\Omega)}.$
Combining this estimate with Young’s inequality, we obtain for all $\zeta_{1}>0$
$\alpha^{-1}N_{h}^{\Delta^{2}}(u_{h},u_{h})=2(\nabla(\Delta u_{h})\cdot\bm{\nu},u_{h})_{L^{2}(\partial\Omega)}\leq C\zeta_{1}h^{-3}\|\Delta u_{h}\|^{2}_{L^{2}(\Omega)}+\zeta_{2}^{-1}\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}.$
The second inequality follows with the same argument using the standard inverse inequality.
Let us introduce the following norm on the discrete space $X_{h}$:
(23) $\|u_{h}\|_{\epsilon}^{2}\coloneqq|u_{h}|^{2}_{H^{2}(\Omega)}+\lvert u_{h}\rvert^{2}_{H^{1}(\Omega)}+\epsilon\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}.$
Notice that in the case of impedance boundary conditions, i.e. $\epsilon=0$, we consider the eigenvalue problem (22) to be posed on the space $\tilde{X}_{h}\cap H^{1}_{0}(\Omega)$, and consequently the $H^{2}$-semi norm does indeed define a norm due to the Poincaré inequality.
Lemma 5.4.
For $\eta_{1},\eta_{2}>0$ sufficiently large, the bilinear form $e_{h}(\cdot,\cdot)$ is uniformly coercive on $X_{h}$, i.e. there exists a constant $\gamma>0$ independent of $h$ s.t.
(24) $\lvert e_{h}(u_{h},u_{h})\rvert\geq\gamma\|u_{h}\|_{\epsilon}^{2}\quad\forall u_{h}\in X_{h}.$
Proof 5.5.
For $u_{h}\in X_{h}$ we estimate with the results from Lemma 5.2:
$\left|\epsilon\left(N_{h}^{\Delta^{2}}(u_{h},u_{h})-N_{h}^{\Delta}(u_{h},u_{h})\right)\right|$
$\geq$ $\epsilon\left(-C^{\Delta^{2}}\zeta_{1}\alpha h^{-3}\|\Delta u_{h}\|_{L^{2}(\Omega)}^{2}-C^{\Delta}\zeta_{2}h^{-1}\|\nabla u_{h}\|^{2}_{L^{2}(\Omega)}-\left(\alpha\zeta_{1}^{-1}+\zeta_{2}^{-1}\right)\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}\right).$
Thus, we have that
$|e_{h}(u_{h},u_{h})|\geq$ $\alpha\left(1-\frac{\epsilon C^{\Delta^{2}}\zeta_{1}}{h^{3}}\right)\|\Delta u_{h}\|^{2}_{L^{2}(\Omega)}+\left(1-\frac{\epsilon C^{\Delta}\zeta_{2}}{h}\right)\|\nabla u_{h}\|^{2}_{L^{2}(\Omega)}$
$+\epsilon\alpha\left(\frac{\eta_{1}}{h^{3}}-\zeta_{1}^{-1}\right)\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}+\epsilon\left(\frac{\eta_{2}}{h}-\zeta_{2}^{-1}\right)\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}$
$\geq$ $\gamma\|u_{h}\|^{2}_{\epsilon},$
where, as usually done when analysing Nitsche’s method, we choose $\zeta_{1}$ and $\zeta_{2}$ sufficiently small and assume that the penalty parameters $\eta_{1}$ and $\eta_{2}$ are chosen sufficiently large.
5.2 Analysis of the discrete problem
In this section, we consider the well-posedness of the discrete problem: find $u_{h}\in X_{h}$ s.t.
(25) $a_{h}(u_{h},v_{h})=(f,v_{h})_{L^{2}(\Omega)}\quad\forall v_{h}\in X_{h},$
where the sesquilinear form $a_{h}(\cdot,\cdot)$ is defined through
(26) $a_{h}(u_{h},v_{h})\coloneqq a(u_{h},v_{h})+\epsilon\left(N^{\Delta^{2}}_{h}(u_{h},v_{h})-N^{\Delta}_{h}(u_{h},v_{h})+S_{h}(u_{h},v_{h})\right).$
Here $a(\cdot,\cdot)$ is the sesquilinear form defined in (2) or (13), $N^{\Delta^{2}}_{h}$, $N^{\Delta}_{h}$ and $S_{h}$ are the Nitsche terms defined in (21) and as before, $\epsilon$ is either equal to one in the sound soft case and or vanishes in the impedance case. Due to the Nitsche terms, $a_{h}(\cdot,\cdot)\not=a(\cdot,\cdot)$ and $X_{h}\not\subset X$ if $\epsilon=1$, so our discretization is non-conforming in the sound soft case. In contrast, the discretization is conforming in the impedance case.
To obtain continuity of $a_{h}(\cdot,\cdot)$ in either case, we introduce the norm
(27) $\|u_{h}\|^{2}_{h,\epsilon}=\|u_{h}\|^{2}_{\epsilon}+\epsilon\left(h^{3}\|\nabla(\Delta u_{h})\|^{2}_{L^{2}(\partial\Omega)}+h\|\nabla u_{h}\|^{2}_{L^{2}(\partial\Omega)}\right).$
Due to the inverse inequalities applied in Lemma 5.2, we have that $\|u_{h}\|^{2}_{\epsilon}\simeq\|u_{h}\|^{2}_{h,\epsilon}$ for all $u_{h}\in X_{h}$ with constants independent of $h$. By construction there exists a constant $C_{\text{cont}}$ such that $a_{h}(u_{h},v_{h})\leq C_{\text{cont}}\|u_{h}\|_{h,\epsilon}\|v_{h}\|_{h,\epsilon}$.
In the following, we denote by $\mathcal{A}_{h}\in\mathcal{L}(X_{h})$ the operator associated with the sesquilinear form $a_{h}(\cdot,\cdot)$. For the impedance case, we recall the definition of $K$
$(Ku_{h},v_{h})_{H^{2}}\coloneqq-\alpha(\Delta u_{h},\nabla v_{h}\cdot\bm{\nu})_{L^{2}(\partial\Omega)}+\alpha i\theta(\Delta u_{h},v_{h})_{L^{2}(\partial\Omega)}-i\theta(u_{h},v_{h})_{L^{2}(\partial\Omega)},$
and set $A_{h}\coloneqq\mathcal{A}_{h}-K\in\mathcal{L}(X_{h})$. For the sound soft case, we have that $K=0$ and set $A_{h}\coloneqq\mathcal{A}_{h}\in\mathcal{L}(X_{h})$.
To show the well-posedness of the discrete problem (25), we will show that $\mathcal{A}_{h}$ is uniformly Th-coercive on $X_{h}$. In particular, it suffices to show that $A_{h}=\mathcal{A}_{h}-K$ is uniformly Th-coercive on $X_{h}$. In the impedance case, this statement is nontrivial and originates from the fact that the compact perturbation $K$ can be neglected asymptotically. The following theorem [^2] Note that this theorem implicitly uses the fact that the sequence of discrete spaces $(X_{h})_{h>0}$ is dense in $X$, i.e. that the approximation property is satisfied. formalizes this intuition and since it is only applied in the impedance case, we state it in a conforming setting. We refer to [23, Thm. 3] for a more general result and to [36, Sec. 8.9] or [4] for similar discussion on compactly perturbed coercive problems.
Theorem 5.6.
Let $\mathcal{A}\in\mathcal{L}(X)$ be injective and weakly T-coercive, that is $\mathcal{A}=A+K$, where $A\in\mathcal{L}(X)$ is T-coercive and $K\in\mathcal{L}(X)$ is compact. If there exists a family of bijective operators $(T_{h})_{h>0}$, $T_{h}\in\mathcal{L}(X_{h})$, such that $A$ is uniformly $T_{h}$-coercive on $X_{h}$, then there exists $h_{0}>0$ such that $\mathcal{A}$ is uniformly $T_{h}$-coercive on $X_{h}$ for all $h\leq h_{0}$.
Proof 5.7.
See e.g. [16, Ex. 26.5] which is based on [43]. For completeness, we provide a proof in Appendix A.
To define a discrete analogue of $T$, let $\{\lambda^{(i)}_{h}\}$ and $\{e^{(i)}_{h}\}$ be the discrete eigenvalues and eigenfunctions obtained from the eigenvalue problem (22). As in the continuous case, we renormalize the eigenfunctions with respect to the $\|\cdot\|_{\epsilon}$-norm. We define the discrete operator $T_{h}\in\mathcal{L}(X_{h})$ as
$W_{h}\coloneqq\text{span}_{0\leq i\leq i_{\ast}}\{e^{(i)}_{h}\},\qquad T_{h}\coloneqq\operatorname{id}_{X_{h}}-2P_{W_{h}},$
where $P_{W_{h}}:X_{h}\to W_{h}$ is the orthogonal projection onto $W_{h}$. As before, we have that $T_{h}^{2}=\operatorname{id}$, thus $T_{h}$ is bijective. Notice that we are assuming that $h$ is small enough such that $\lambda^{(i_{\ast})}_{h}<k^{2}$.
Theorem 5.8.
For $h$ small enough such that $\lambda^{(i_{\ast})}_{h}<k^{2}$ and $\eta_{1},\eta_{2}>0$ sufficiently large, the operator $A_{h}\in\mathcal{L}(X_{h})$ is uniformly Th-coercive on $X_{h}$.
Proof 5.9.
For any $u_{h}\in X_{h}$, we have that due to the smallness assumption on $h$
$(A_{h}T_{h}u_{h},u_{h})_{H^{2}}$ $=\sum_{0\leq i\leq i_{\ast}}C(\lambda^{(i)}_{h})(k^{2}-\lambda^{(i)}_{h})(u_{h}^{(i)})^{2}+\sum_{i>i_{\ast}}C(\lambda^{(i)}_{h})(\lambda^{(i)}_{h}-k^{2})(u_{h}^{(i)})^{2}$
$\geq C(\epsilon)\gamma\|u_{h}\|^{2}_{\epsilon},$
where $\gamma\coloneqq\min_{i\in\mathbb{N}}C(\lambda^{(i)}_{h})\{|\lambda^{(i)}_{h}-k^{2}|\}>0$ is uniform in $h$, with $C(\lambda^{(i)}_{h})$ being the renormalization constant. Thus $A_{h}$ is uniformly Th-coercive on $X_{h}$.
For sound soft boundary conditions, this immediately shows that the discrete problem is well-posed, since $\mathcal{A}_{h}=A_{h}$. For impedance boundary conditions, we can apply Theorem 5.6 to infer the asymptotic stability of $\mathcal{A}_{h}$ from the stability of $A_{h}$.
Corollary 5.10.
The discrete problem is consistent, i.e. for any $u\in H^{2}(\Omega)$ solving (6), we have that
$a_{h}(u-u_{h},v_{h})=0\quad\forall v_{h}\in X_{h}.$
Proof 5.11.
For $\epsilon=0$, the statement is trivial since $a_{h}(\cdot,\cdot)=a(\cdot,\cdot)$. For $\epsilon=1$, the statement follows from partial integration as usual for the Nitsche’s method.
Thus, we have the following best approximation result.
Theorem 5.12.
Suppose that the assumptions of Theorem 5.8 hold. Let $u\in X$ be the solution of (6) and $u_{h}\in X_{h}$ be the solution of the discrete problem (25). Then, there exists a constant $C>0$ such that
$\|u-u_{h}\|_{h,\epsilon}\leq C\inf_{v_{h}\in X_{h}}\|u-v_{h}\|_{h,\epsilon}.$
Proof 5.13.
For $v_{h}\in X_{h}$ the triangle inequality yields that
$\|u-u_{h}\|_{h,\epsilon}\leq\|u-v_{h}\|_{h,\epsilon}+\|v_{h}-u_{h}\|_{h,\epsilon}.$
Using the uniform Th-coercivity of $a_{h}(\cdot,\cdot)$ and Corollary 5.10, we obtain for the second term
$\gamma\|u_{h}-v_{h}\|_{h,\epsilon}^{2}$ $\leq a_{h}(T_{h}(u_{h}-v_{h}),u_{h}-v_{h})\leq\|T_{h}\|_{\mathcal{L}(X_{h})}a_{h}(u-v_{h},u_{h}-v_{h})$
$\leq C_{\text{cont}}\|T_{h}\|_{\mathcal{L}(X_{h})}\|u-v_{h}\|_{h,\epsilon}\|u_{h}-v_{h}\|_{h,\epsilon}.$
Dividing by $\|u_{h}-v_{h}\|_{h,\epsilon}$ and taking the infimum over $v_{h}\in X_{h}$ we obtain
$\|u-u_{h}\|_{h,\epsilon}\leq\frac{C_{\text{cont}}\|T_{h}\|_{\mathcal{L}(X_{h})}}{\gamma}\inf_{v_{h}\in X_{h}}\|u-v_{h}\|_{h,\epsilon}.$
Remark 5.14 (Sound hard boundary conditions).
The case of sound hard boundary conditions can be analyzed similarly. To implement the boundary conditions with Nitsche’s method, we set $N^{\Delta^{2}}_{h}=0$ and redefine the penalization term as
$S_{h}(u_{h},v_{h})\coloneqq\frac{\eta}{h}(\partial_{n}u_{h},\partial_{n}v_{h})_{L^{2}(\partial\Omega)}.$
The analysis requires the study of a different discrete and continuous eigenvalue problem with respect to (16) and (22). This analysis is similar and we omit the details.
6 Discretization of the nematic Helmholtz–Korteweg equation ($\beta>0$)
We extend the discrete analysis from the previous section to the case where $\beta>0$. In addition to the Nitsche terms defined by (21), we define
$N_{h}^{\beta}(u_{h},v_{h})$ $\coloneqq\beta(\nabla(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n})\cdot\bm{\nu},v_{h})_{L^{2}(\partial\Omega)}+\beta(u_{h},\nabla(\mathbf{n}^{T}(\mathcal{H}v_{h})\mathbf{n})\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$
$S_{h}^{\beta}(u_{h},v_{h})$ $\coloneqq\frac{\beta\eta_{3}}{h^{3}}(u_{h},v_{h})_{L^{2}(\partial\Omega)}.$
To adjust the T-coercivity argument on the discrete level, we consider the following discrete eigenvalue problem: find $u_{h}\in\tilde{X}_{h}\subseteq X_{h}$, $\lambda\in\mathbb{R}$ such that
(28) $e_{h}(u_{h},v_{h})=\lambda(u_{h},v_{h})_{L^{2}(\Omega)}\quad\forall v_{h}\in\tilde{X}_{h}\subseteq X_{h},$
where the bilinear form $e_{h}(\cdot,\cdot)$ is defined through
$e_{h}(u_{h},v_{h})$ $\coloneqq\alpha(\Delta u_{h},\Delta v_{h})_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n},\Delta v_{h})_{L^{2}(\Omega)}+(\nabla u_{h},\nabla v_{h})_{L^{2}(\Omega)}$
$+\epsilon\left(N^{\Delta^{2}}_{h}(u_{h},v_{h})-N^{\Delta}_{h}(u_{h},v_{h})+N^{\beta}_{h}(u_{h},v_{h})+S_{h}(u_{h},v_{h})+S^{\beta}_{h}(u_{h},v_{h})\right).$
As in the previous section, we set $\epsilon=1$ for the sound soft case and $\epsilon=0$ for the impedance case, where we assume for the latter case that the boundary conditions $u_{h}=0$ and $\Delta u_{h}=0$ on $\partial\Omega$ are imposed through the discrete space $\tilde{X}_{h}$ and in a natural way, respectively. For the former case, we have that $\tilde{X}_{h}=X_{h}$ and the boundary conditions are imposed through Nitsche’s method.
6.1 Analysis of the discrete eigenvalue problem
Lemma 6.1.
There exists a constant $C>0$ such that the following inverse trace inequality holds
(29) $\|\nabla(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n})\cdot\bm{\nu}\|^{2}_{L^{2}(\partial\Omega)}\leq Ch^{-3}|u_{h}|^{2}_{H^{2}(\Omega)}.$
Proof 6.2.
We set $\mathcal{K}_{h}\coloneqq\{v_{h}\in X_{h}:\nabla(\mathbf{n}^{\top}\mathcal{H}v_{h}\mathbf{n})=0\}$ such that $X_{h}=\mathcal{K}_{h}^{\perp}\oplus\mathcal{K}_{h}$ and consider the following eigenvalue problem on $\mathcal{K}_{h}^{\perp}$: find $u_{h}\in\mathcal{K}_{h}^{\perp}$, $\tilde{\lambda}_{h}\in\mathbb{R}$ such that for all $v_{h}\in\mathcal{K}_{h}^{\perp}$ we have
(30) $h(\nabla(\mathbf{n}^{\top}\mathcal{H}u_{h}\mathbf{n})\cdot\bm{\nu},\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n})\cdot\bm{\nu})_{L^{2}(\partial\Omega)}=\tilde{\lambda}_{h}(\nabla(\mathbf{n}^{T}\mathcal{H}u_{h}\mathbf{n}),\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n}))_{L^{2}(\Omega)}.$
The left hand-side is bounded and the right hand-side is coercive $\mathcal{K}_{h}^{\perp}$, so the problem is well-posed, and the associated eigenvalues are positive and finite. Further, the min-max characterization yields for the maximal eigenvalue $\tilde{\lambda}_{h,\max}$ that
(31) $\tilde{\lambda}_{h,\max}=\sup_{v_{h}\in\mathcal{K}_{h}^{\perp}\setminus\{0\}}\frac{h(\nabla(\mathbf{n}^{\top}\mathcal{H}v_{h}\mathbf{n})\cdot\bm{\nu},\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n})\cdot\bm{\nu})_{L^{2}(\partial\Omega)}}{(\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n}),\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n}))_{L^{2}(\Omega)}}.$
Thus, we obtain that $\|\nabla(\mathbf{n}^{\top}\mathcal{H}u_{h}\mathbf{n})\cdot\bm{\nu}\|_{L^{2}(\partial\Omega)}^{2}\leq Ch^{-1}\|\nabla(\mathbf{n}^{T}\mathcal{H}u_{h}\mathbf{n}^{T})\|_{L^{2}(\Omega)}^{2}$. Then, we use the a standard inverse inequality and that $\|\mathbf{n}\|^{2}_{L^{\infty}}=1$ to estimate
$h^{-1}\|\nabla(\mathbf{n}^{\top}\mathcal{H}u_{h}\mathbf{n}^{T})\|_{L^{2}(\Omega)}^{2}\leq Ch^{-3}\|\mathcal{H}u_{h}\|_{L^{2}(\Omega)}^{2}\leq Ch^{-3}|u_{h}|_{H^{2}(\Omega)}^{2}.$
With the result from Lemma 6.1 we can follow the same argument as in the previous section to show the uniform coercivity of $e_{h}(\cdot,\cdot)$ with respect to the norm defined in (23) and conclude the well-posedness of the discrete eigenvalue problem.
Lemma 6.3.
The bilinear form $e_{h}(\cdot,\cdot)$ is uniformly coercive on $\tilde{X}_{h}$.
Proof 6.4.
In addition to the estimates from Lemma 5.4, we use the inverse trace inequality from Lemma 6.1 to estimate with Young’s inequality for $\zeta_{3}>0$
(32) $\left|\epsilon N^{\beta}_{h}(u_{h},u_{h})\right|\geq-\epsilon\beta\left(C^{\beta}\zeta_{3}|u_{h}|^{2}_{H^{2}(\Omega)}-\zeta_{3}^{-1}\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}\right).$
Consequently, we obtain
$|e_{h}(u_{h},u_{h})|\geq$ $\alpha\left(1-\frac{\epsilon C^{\Delta^{2}}\zeta_{1}}{h^{3}}\right)\|\Delta u_{h}\|^{2}_{L^{2}(\Omega)}+\left(1-\frac{\epsilon C^{\Delta}\zeta_{2}}{h}\right)\|\nabla u_{h}\|^{2}_{L^{2}(\Omega)}$
$-\epsilon\beta C^{\beta}\zeta_{3}|u_{h}|^{2}_{H^{2}(\Omega)}+\epsilon\beta\left(\frac{\eta_{3}}{h^{3}}-\zeta_{3}^{-1}\right)\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}$
$+\epsilon\alpha\left(\frac{\eta_{1}}{h^{3}}-\zeta_{1}^{-1}\right)\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}+\epsilon\left(\frac{\eta_{2}}{h}-\zeta_{2}^{-1}\right)\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}$
$\geq$ $\,\gamma\|u_{h}\|^{2}_{\epsilon},$
where we choose $\zeta_{1},\zeta_{2},\zeta_{3}$ sufficiently small and assume that the penalty parameters $\eta_{1},\eta_{2},\eta_{3}$ are chosen sufficiently large.
6.2 Analysis of the discrete problem
We consider the stability of the discrete problem: find $u_{h}\in X_{h}$ such that
(33) $a_{h}(u_{h},v_{h})=(f,v_{h})_{L^{2}(\Omega)}\quad\forall v_{h}\in X_{h},$
where the sesquilinear form $a_{h}(\cdot,\cdot)$ is defined through
(34) $a_{h}(u_{h},v_{h})\coloneqq a(u_{h},v_{h})+\epsilon\big{(}$ $N^{\Delta^{2}}_{h}(u_{h},v_{h})-N^{\Delta}_{h}(u_{h},v_{h})$
$+N^{\beta}_{h}(u_{h},v_{h})+S_{h}(u_{h},v_{h})+S^{\beta}_{h}(u_{h},v_{h})\big{)}.$
To account for the nematic terms, we modify the continuity norm (27) such that the sesquilinear form $a_{h}(\cdot,\cdot)$ is continuous on $X_{h}$:
(35) $\|u_{h}\|^{2}_{h,\epsilon}\coloneqq\|u_{h}\|^{2}_{\epsilon}+\epsilon\left(h^{3}(\|\nabla(\Delta u_{h})\|^{2}_{L^{2}(\partial\Omega)}+\|\nabla(\mathbf{n}^{\top}\mathcal{H}u_{h}\mathbf{n})\|_{L^{2}(\partial\Omega)}^{2})+h\|\nabla u_{h}\|^{2}_{L^{2}(\partial\Omega)}\right).$
As before, we denote by $\mathcal{A}_{h}\in\mathcal{L}(X_{h})$ the linear operator induced by the bilinear form $a_{h}(\cdot,\cdot)$ and set $A_{h}=\mathcal{A}_{h}-K$ where $K\in\mathcal{L}(X_{h})$ is defined as
$(Ku_{h},v_{h})_{H^{2}}\coloneqq$ $-\alpha(\Delta u_{h},\nabla v_{h}\cdot\bm{\nu})_{L^{2}(\partial\Omega)}+\alpha i\theta(\Delta u_{h},v_{h})_{L^{2}(\partial\Omega)}-i\theta(u_{h},v_{h})_{L^{2}(\partial\Omega)}$
$+\beta i\theta(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n},v_{h})_{L^{2}(\partial\Omega)}-\beta(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n},\nabla v_{h}\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$
in the impedance case and $K=0$ in the sound soft case. Due to Theorem 5.6, it suffices to show that the operator $A_{h}$ is uniformly Th-coercive to conclude the well-posedness of both cases.
Let $\{\lambda^{(i)}_{h}\}$ and $\{e^{(i)}_{h}\}$ be the discrete eigenvalues and eigenfunctions obtained from the eigenvalue problem (28), where the latter are renormalized with respect to the $\|\cdot\|_{\epsilon}$-norm. Then, we define
$W_{h}\coloneqq\text{span}_{0\leq i\leq i_{\ast}}\{e^{(i)}_{h}\},\qquad T_{h}\coloneqq\operatorname{id}_{X_{h}}-2P_{W_{h}},$
where $P_{W_{h}}:X_{h}\to W_{h}$ is the orthogonal projection onto $W_{h}$. As before, we have that $T_{h}^{2}=\operatorname{id}$, thus $T_{h}\in\mathcal{L}(X_{h})$ is bijective.
Theorem 6.5.
For $h$ small enough such that $\lambda_{h}^{(i_{\ast})}<k^{2}$, $\beta$ sufficiently small and $\eta_{1},\eta_{2}>0$ large enough, the operator $A_{h}\in L(X_{h})$ is uniformly $T_{h}$-coercive on $X_{h}$.
Proof 6.6.
Due to the modified eigenvalue problem (28), the proof follows with exactly the same argument as in the proof of Theorem 5.8.
Furthermore, we can also use the same argument as in Theorem 5.12 to obtain the following best approximation result.
Theorem 6.7.
Let the assumptions of Theorem 6.5 be satisfied. For $u\in X$ solving (6) and $u_{h}\in X_{h}$ solving the discrete problem (33), there exists a constant $C>0$ such that
$\|u-u_{h}\|_{h,\epsilon}\leq C\inf_{v_{h}\in X_{h}}\|u-v_{h}\|_{h,\epsilon}.$
Remark 6.8 (Ensuring quasi-optimaltiy).
To ensure that the smallness assumption on $h$ is satisfied, we can adopt a similar scheme as introduced in [41] for the Helmholtz problem. By solving the discrete eigenvalue problem (22) on a sequence of refined meshes, the criterion $\lambda^{(i_{\ast})}_{h}<k^{2}$ can be checked. If the criterion is met, the assumptions of Theorem 5.8 and Theorem 6.5 are satisfied, otherwise the mesh is refined and the procedure is repeated. To reduce the required number of degrees of freedom, an adaptive error estimator for the first $i_{\ast}$ eigenfunctions can be used.
7 Numerical examples
The analysis above is abstract and applies to any $C^{1}$ conforming finite element space, including the Argyris element [2] ($p\geq 5$), the Hsieh–Clough–Tocher macroelement defined on the Alfeld split of a triangular mesh [13] ($p\geq 3$), and the Morgan–Scott element [33, 1] ($p=5$). Here we will present numerical simulations implemented using the Firedrake finite element library and ngsPETSc [24, 8, 5]. We test the Argyris element and the Hsieh–Clough–Tocher macroelements.
7.1 Quasi optimal convergence
Plane waves are solutions of the Helmholtz equation that remain constant over a plane perpendicular to the direction of propagation. They are a natural choice for the numerical validation of the Helmholtz–Korteweg equation, since they are also solutions of the latter, i.e.
(36) $u(\mathbf{x})=e^{i(\mathbf{d}\cdot\mathbf{x})},\qquad\mathbf{d}\in\mathbb{C}.$
Substituting (36) into (5), it is easy to observe that plane waves are also exact solution of the nematic Helmholtz–Korteweg equation, under appropriate choice of the wave vector $\mathbf{d}$ [18]. In particular, if $d$ satisfies the following dispersion relation (37), then the plane wave is also a solution of the Helmholtz–Korteweg equation:
(37) $\alpha d^{4}+\beta d^{2}(\mathbf{d}^{\top}\mathbf{n})^{2}+d^{2}-k^{2}.$
In Figure 1 we present the convergence of the $H^{2}$-norm of the error for the Helmholtz–Korteweg equation with data constructed in such a way that the exact solution is a plane wave. We observe that from Theorem 5.12 the error should converge with a rate only determined by the approximation properties of the finite element space. In particular, it is well known, see for example [7, Section 5.9], that one can construct an interpolation operator from $H^{5}(\Omega)$ to the Argyris finite element space that converges with rate $h^{3}$ with respect to the $H^{2}$-norm [2], i.e.
(38) $\|u-I_{h}u\|_{H^{2}(\Omega)}\leq Ch^{4}\|u\|_{H^{5}(\Omega)}.$
Thus, combining (38) with Theorem 5.12 we expect the error to converge with rate $h^{4}$ with respect to the $H^{2}$-norm, as confirmed by Figure 1. Similarly, if we consider the Hsieh–Clough–Tocher macroelement [13] we expect the error to converge with rate $h^{2}$ with respect to the $H^{2}$-norm, as confirmed by Figure 1.
[…]
Figure 1: The convergence of the $H^{2}$-norm of the error for the Helmholtz–Korteweg equation for different values of $k$ (top row) and the corresponding manufactured solution (bottom row).
7.2 Anisotropic Gaussian pulse
We study the anisotropic effect of the nematic term $\beta\nabla\cdot\nabla\left(\mathbf{n}^{\top}\mathcal{H}u\mathbf{n}\right)$ on the propagation of a symmetric Gaussian pulse of the form
(39) $f(x,y)=\exp\left(-(40^{2})\left[\left(x-\frac{1}{2}\right)^{2}+\left(y-\frac{1}{2}\right)^{2}\right]\right).$
Several authors have studied the propagation of plane waves in the context of the nematic Helmholtz–Korteweg equation [18, 42, 37]. In particular, it has been shown that the nematic term causes an anisotropic speed of propagation of plane waves, which is greater in the direction of the nematic director $\mathbf{n}$. In Figure 2 we present numerical simulations of the propagation of a symmetric Gaussian pulse by the nematic Helmholtz–Korteweg equation with sound soft boundary conditions, and in Figure 3 we impose impedance boundary conditions. Both figures show that the anisotropic speed of propagation observed for plane waves is also present for the Gaussian pulse, independent of the choice of boundary conditions.
Figure 2: The propagation of a symmetric Gaussian pulse by the nematic Helmholtz–Korteweg equation with sound soft boundary conditions when $\beta=0$ (first and left-most figure) and when $\beta=5\cdot 10^{-3}$ and $\mathbf{n}$ is aligned with the $x$-axis (second figure), the diagonal (third figure) and the $y$-axis (last and right-most figure). The parameters here are $k=40$, $\alpha=10^{-2}$.
Figure 3: The propagation of a symmetric Gaussian pulse by the nematic Helmholtz–Korteweg equation with impedance boundary conditions when $\beta=0$ (first and left-most figure) and when $\beta=5\cdot 10^{-3}$ and $\mathbf{n}$ is aligned with the $x$-axis (second figure), the diagonal (third figure) and the $y$-axis (last and right-most figure). The parameters here are $k=40$, $\alpha=10^{-2}$.
7.3 The Mullen–Lüthi-Stephen experiment
The Mullen–Lüthi-Stephen experiment [34] consists of a planar acoustic wave propagating in a nematic liquid crystal, where the nematic director $\mathbf{n}$ kept fixed in the $x$-direction via a magnetic field. The experiment reveals that the speed of propagation of the acoustic wave is anisotropic, being greater in the direction of the nematic director $\mathbf{n}$. We consider a variant of the Mullen–Lüthi-Stephen experiment, where the nematic director $\mathbf{n}$ is kept fixed in the $x$-direction in the central region of the domain and in the $y$-direction in the outer region of the domain. We then study the propagation of a planar acoustic wave coming from the top of the domain. We can clearly see from Figure 4 that the speed of propagation of the acoustic wave is anisotropic, being greater in the central region of the domain where the nematic director $\mathbf{n}$ is aligned with the direction of propagation of the acoustic wave. Furthermore, we can observe that the damping along the Robin boundary conditions imposed on the sides of the domain depends on the orientation of the nematic director $\mathbf{n}$. In particular, the damping is greater in the region where the nematic director $\mathbf{n}$ is aligned with the direction of propagation of the acoustic wave. This type of phenomena has been predicted from the analysis of the nematic Helmholtz–Korteweg equation [18]. Lastly we would like to remark that discontinuity in the nematic director field presented in Figure 4 cannot be achieved if the nematic director field is a general solution of the Euler–Lagrange equations associated with the Oseen–Frank energy density, and should be regarded as an approximation of a solution of the Oseen–Frank model exhibiting a sharp transition between two constant orientations.
Figure 4: The anisotropic propagation of a planar acoustic wave in the modified Mullen–Lüthi-Stephen experiment (top) and the corresponding nematic director $\mathbf{n}$ (bottom). The parameters are $k=40$, $\alpha=10^{-4}$ and $\beta=5\cdot 10^{-5}$.
Acknowledgements
The authors would like to thank Christoph Lehrenfeld and Ilaria Perugia for their valuable comments and suggestions.
Appendix A Proof of Theorem 5.6
Proof A.1.
We follow the argument in [16, Ex. 26.5] which is based on [43] and make slight adaptations for the setting of T-coercivity. Let $R_{h}:V\rightarrow V_{h}$ be the discrete solution map for $A$, i.e. for all $v\in V$ it holds that $(A(R_{h}(v)-v),w_{h})_{V}=0$ for all $w_{h}\in X_{h}$. Since $A$ is uniformly Th-coercive, we have that
$\alpha_{0}\|R_{h}(v)\|_{X}$ $\leq\frac{|(AR_{h}(v),T_{h}R_{h}(v))_{X}|}{\|R_{h}(v)\|_{X}}=\frac{|(Av,T_{h}R_{h}(v))_{X}|}{\|R_{h}(v)\|_{X}}$
$\leq\|A\|_{\mathcal{L}(X,X)}\|T_{h}\|_{\mathcal{L}(X,X)}\|v\|_{X}.$
Thus, it holds that $\|R_{h}\|_{\mathcal{L}(X,X)}\leq(\|A\|_{\mathcal{L}(X,X)}\|T_{h}\|_{\mathcal{L}(X,X)}/\alpha_{0})$ and we conclude that
$\|R_{h}(v)-v\|_{X}\leq\left(1+\frac{\|A\|_{\mathcal{L}(X,X)}\|T_{h}\|_{\mathcal{L}(X,X)}}{\alpha_{0}}\right)\inf_{w_{h}\in X_{h}}\|v-w_{h}\|_{X}.$
Therefore, we have that $R_{h}\rightarrow\operatorname{Id}_{X}$ pointwise as $h\rightarrow 0$ which implies that $R_{h}\rightarrow\operatorname{Id}_{X}$ uniformly as $h\rightarrow 0$ on compact sets [16, Rem. C.5]. Define $L\coloneqq\operatorname{Id}_{X}+A^{-1}K$ and $L_{h}\coloneqq\operatorname{Id}_{X}+R_{h}A^{-1}K$. Then, we have that $L_{h}\rightarrow L$ in $\mathcal{L}(X,X)$ due to the uniform convergence of $R_{h}$ to $\operatorname{Id}_{X}$ on compact sets and the compactness of $A^{-1}K$. We note that $L$ is bijective and $L^{-1}L_{h}=\operatorname{Id}_{X}-L^{-1}(L-L_{h})$. Taking $h<h_{1}$ small enough such that $\|L^{-1}(L-L_{h})\|_{\mathcal{L}(X,X)}\leq 1/2$ ensures that $L_{h}$ is invertible with
(40) $L_{h}^{-1}L=\sum_{k\in\mathbb{N}}\left(L^{-1}(L-L_{h})\right)^{k},$
such that $\|L_{h}^{-1}\|_{\mathcal{L}(X,X)}\leq 2\|L^{-1}\|_{\mathcal{L}(X,X)}\eqqcolon C_{L}$. Consequently, we have that
$\frac{\alpha_{0}}{C_{L}}\|v_{h}\|^{2}_{X}$ $\leq\alpha_{0}\|L_{h}(v_{h})\|^{2}_{X}\leq|(AL_{h}(v_{h}),T_{h}v_{h})_{X}|$
$\leq|(A(L-L_{h})(v_{h}),T_{h}v_{h})_{X}|+|(\mathcal{A}v_{h},T_{h}v_{h})_{X}|$
$\leq\|A\|_{\mathcal{L}(X,X)}\|T_{h}\|_{\mathcal{L}(X,X)}\|L-L_{h}\|_{\mathcal{L}(X,X)}\|v_{h}\|_{X}^{2}+|(\mathcal{A}v_{h},T_{h}v_{h})_{X}|,$
where the second to last step follows from $AL=A(\operatorname{Id}_{X}+A^{-1}K)=\mathcal{A}$. For $h<h_{2}$ small enough, the first term on the right-hand side can be bounded by $\alpha_{0}/2C_{L}$ such that for $h<h_{0}=:\min(h_{1},h_{2})$ it holds that
(41) $\frac{\alpha_{0}}{2C_{L}}\|v_{h}\|^{2}_{X}\leq|(\mathcal{A}v_{h},T_{h}v_{h})_{X}|.$
Thus, $\mathcal{A}$ is uniformly Th-coercive.