### v2

# Analysis and numerical analysis of the nematic Helmholtz–Korteweg equation

Abstract
We analyse the nematic Helmholtz–Korteweg equation, a variant of the classical Helmholtz equation that describes time-harmonic wave propagation in calamitic fluids in the presence of nematic order. A prominent example is given by nematic liquid crystals, which can be modeled as nematic Korteweg fluids—that is, fluids whose stress tensor depends on density gradients and on a nematic director describing the orientation of the anisotropic molecules. These materials exhibit anisotropic acoustic properties that can be tuned by external electromagnetic fields, making them attractive for potential applications such as tunable acoustic resonators. We prove the existence and uniqueness of solutions to this equation in two and three dimensions for suitable (nonresonant) wave numbers and propose a convergent discretisation for its numerical solution. The discretisation of this problem is nontrivial as it demands high regularity and involves unfamiliar boundary conditions; we address these challenges by using high-order conforming finite elements and enforcing the boundary conditions with Nitsche’s method. We illustrate our analysis with numerical simulations in two dimensions.
keywords:
Korteweg fluids, nematic liquid crystals, Helmholtz equation, wave propagation
1 Introduction
In this work we consider the analysis and numerical analysis of the nematic Helmholtz–Korteweg equation
$\alpha\Delta^{2}u+\beta\nabla\cdot\nabla\left(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\right)-\Delta u-k^{2}u=f,$ (1)
where $u$ represents the density perturbation of a calamitic fluid (a fluid composed of rodlike molecules) in the presence of nematic ordering (where the molecules locally align, like arrows in a quiver), $\mathcal{H}u$ is the Hessian matrix of $u$, $\mathbf{n}$ is a piecewise constant unit-vector field describing the local average orientation of the molecules in the fluid, and the parameters $\alpha\gg\beta\geq 0$ are material constitutive parameters known as acoustic susceptibilities [VirgaNematoacoustics].
The most prominent example of such fluids are nematic liquid crystals. These fluids exhibit long-range orientational order within particular temperature or concentration ranges. Molecules close to each other tend to align their molecular axes, giving rise to a macroscopic ordering. Nematic liquid crystals have found widespread application in display technology and the design of novel devices due to their anisotropic acoustic, elastic, and optical properties [deGennes]. Furthermore, a remarkable feature of nematic liquid crystals is that these properties can be tuned by external electromagnetic fields.
The nematic Helmholtz–Korteweg equation, recently derived by a subset of the authors [farrell2024b], is a novel partial differential equation modeling time-harmonic acoustic wave propagation in nematic liquid crystals. Within this model, nematic liquid crystals are treated as Korteweg fluids with an additional anisotropic term in the Cauchy stress tensor $\sigma$ depending on the nematic director $\mathbf{n}$ [VirgaNematoacoustics], i.e.
$\sigma=-pI-\alpha u\left(\nabla u\otimes\nabla u\right)-\beta\left(\nabla u\cdot\mathbf{n}\right)\nabla u\otimes\mathbf{n},$ (2)
where $p$ is the fluid pressure. The term with $\alpha$ is the Korteweg term, accounting for stresses caused by spatial variations in density [Giovangigli], while the term with $\beta$ is the anisotropic term arising due to nematic ordering. The solutions of the nematic Helmholtz–Korteweg equation exhibit surprising phenomena not permitted by the classical Helmholtz equation (recovered by setting $\alpha=\beta=0$), including anisotropy in the propagation of sound, penetration depth, and scattering [farrell2024b].
Let $\Omega\subset\mathbb{R}^{d}$, $d\in\{2,3\}$ is be a bounded convex domain with smooth boundary and let $\bm{\nu}$ be the outward-facing unit normal on $\partial\Omega$. For a given source term $f\in L^{2}(\Omega)$, we consider the boundary value problem
$\alpha\Delta^{2}u+\beta\nabla\cdot\nabla\left(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\right)-\Delta u-k^{2}u$ $=f$ $\text{ in }\Omega,$ (3)
$\mathcal{B}u$ $=0$ $\text{ on }\partial\Omega,$
where $\mathcal{B}$ is an operator encoding the boundary conditions. In particular, $\mathcal{B}$ can encode sound soft, sound hard, or impedance boundary conditions.
Sound soft boundary conditions impose that the acoustic pressure vanishes on the boundary, which corresponds to
$\mathcal{B}u=(u,\alpha\Delta u+\beta\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}).$ (4)
Sound hard boundary conditions impose that the normal component of the fluid velocity vanishes on the boundary, which corresponds to
$\mathcal{B}u=(\nabla u\cdot\bm{\nu},\alpha\nabla(\Delta u)\cdot\bm{\nu}+\beta\nabla(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n})\cdot\bm{\nu}).$ (5)
Lastly, impedance boundary conditions impose that the normal component of the fluid velocity is proportional to the fluid velocity, which corresponds to
$\mathcal{B}u=(\nabla u\cdot\bm{\nu}-i\theta u,\alpha\nabla(\Delta u)\cdot\bm{\nu}-i\theta(\alpha\Delta u+\beta\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}-\beta\nabla(\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n})\cdot\bm{\nu})),$ (6)
for specified impedance $\theta$.
For suitable (nonresonant) wave numbers $k$, we prove existence and uniqueness of solutions to (3) for each type of boundary conditions introduced above, see Theorem 3.4. We then analyse a classical $H^{2}$-conforming finite element discretisation of these problems, yielding the first convergent numerical method for the nematic Helmholtz–Korteweg equation. Compared to the analysis and discretisation of singularly perturbed biharmonic problems, the nematic Helmholtz–Korteweg equation presents additional challenges due to the anisotropic Korteweg term and the nonstandard boundary conditions introduced in [farrell2024b]. To tackle these challenges, we use a weak T-coercivity technique [BCS02, DCZ10, Halla21], where we treat the boundary terms as compact perturbations and analyse the underlying eigenvalue problem to establish T-coercivity. On the discrete level, we use Nitsche’s method [Nit71] to enforce the boundary conditions in the sound soft case.
We apply the proposed numerical method to two-dimensional problems, illustrating the capabilities of the method and verifying that the nematic Helmholtz–Korteweg equation exhibits anisotropic wave propagation [farrell2024b, VirgaNematoacoustics, MullenLuti].
In particular, we show that nematic Helmholtz–Korteweg equation correctly reproduces the experiment where anisotropic propagation speed was first observed (known as the Mullen–Lüthi–Stephen experiment [MullenLuti]), thus partially validating the idea first proposed in [VirgaNematoacoustics] and further studied in [farrell2024b]. Lastly, the nematic Helmholtz–Korteweg equation is used to perform numerical simulations of acoustic resonance in a cavity filled with a nematic liquid crystal, paving the way to simulation-aided design of new acoustic devices based on nematic liquid crystals.
We remark that the nematic Helmholtz–Korteweg equation is a valid model for the acoustic propagation in nematic calamitic fluids provided that the nematic director field $\mathbf{n}$ is undistorted, i.e. $\nabla\mathbf{n}\equiv 0$. This assumption is valid at small length scale; for larger length scales one must also consider the elastic effects related to distortions of the nematic director field.
2 Notation & weak formulation
We denote by $X$ a Hilbert space with associated inner product $(\cdot,\cdot)_{X}$ and by $\mathcal{L}(X)$ the space of bounded linear operators acting on $X$. For a given bounded sesquilinear form $a(\cdot,\cdot)$, we denote by $A\in\mathcal{A}$ the associated linear operator. To analyse the weak formulation of (1) introduced below, we use the concept of (weak) T-coercivity [BCS02, DCZ10], which was introduced to treat sign shifting coefficients [DCZ10, BCC14, BCC18, Halla21]. It can also be used as an equivalent alternative to the inf-sup condition in a general setting, for example for the analysis of Helmholtz-like problems [Cia12, vBZ24, buffaPerugia, BoffiJEE02], Galbrun’s equation [HH21, HLS22, Halla23, vB23, HLvB25], or the Stokes problem [EJ25] and their respective finite element discretisations.
Definition 2.1.
We call a sesquilinear form $a:X\times X\to\mathbb{C}$ T-coercive if there exists a bijective operator $T\in\mathcal{L}(X)$ and a constant $\gamma>0$ such that
$\Re\{a(Tu,u)\}\geq\gamma\|u\|^{2}_{X}\quad\forall u\in X.$ (7)
Equivalently, we require that $AT$ is coercive, where $A\in\mathcal{L}(X)$ is the operator associated to the sesquilinear form $a(\cdot,\cdot)$.
It can be shown that T-coercivity is equivalent to the inf-sup condition [Cia12] and thus, it is a necessary and sufficient condition for the well-posedness of the problem.
Definition 2.2 (weak T-coercivity).
An operator $A\in\mathcal{L}(X)$ is called weakly T-coercive if there exists a bijective operator $T\in\mathcal{L}(X)$ and a compact operator $K\in\mathcal{L}(X)$ such that $AT+K$ is bijective.
It is well-known that weakly T-coercive operators are Fredholm with index zero and therefore bijective if and only if they are injective [Halla21, vB23].
To derive the weak formulation of (1), we multiply with a test function $v\in X$ and integrate by parts twice. For the first term (biharmonic), we obtain
$\alpha$ $\int_{\Omega}\nabla\cdot\nabla(\Delta u)v\,\mathrm{d}x=\alpha\int_{\Omega}\Delta u\Delta v\,\mathrm{d}x-\alpha\int_{\partial\Omega}\Delta u\nabla v\cdot\bm{\nu}\,\mathrm{d}s+\alpha\int_{\partial\Omega}\nabla(\Delta u)\cdot\bm{\nu}v\,\mathrm{d}s,$ (8)
and for the second (nematic) term, integration by parts twice yields
$\beta$ $\int_{\Omega}\nabla\cdot\nabla(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n})v\,\mathrm{d}x$
$=\beta\!\int_{\Omega}\!\!\left(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\right)\Delta v\,\mathrm{d}x-\beta\!\int_{\partial\Omega}\!\!\left(\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n}\right)\nabla v\cdot\bm{\nu}\,\mathrm{d}s+\beta\!\int_{\partial\Omega}\nabla\left(\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n}\right)\cdot\bm{\nu}v\,\mathrm{d}s,$ (9)
For the third (Laplacian) term this procedure is standard. The weak form of (3) is given by
$\text{find }u\in X\text{ such that }a(u,v)=(f,v)_{L^{2}(\Omega)}\quad\forall v\in X,$ (10)
where we specify the bilinear form $a(\cdot,\cdot)$ and the space $X$ for each type of boundary condition. For sound soft boundary conditions, we set $X\coloneqq H^{2}(\Omega)\cap H^{1}_{0}(\Omega)$ and define the bilinear form $a(\cdot,\cdot)$ to be
$a(u,v)\coloneqq$ $\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}\!-\!k^{2}(u,v)_{L^{2}(\Omega)},$ (11)
where the boundary terms vanish due to the boundary conditions. For sound hard boundary conditions, we set $X\coloneqq H^{2}(\Omega)$ and define the bilinear form $a(\cdot,\cdot)$ to be
$a(u,v)\coloneqq$ $\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}-k^{2}(u,v)_{L^{2}(\Omega)}$ (12)
- $\alpha(\Delta u,\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$
where the absent boundary terms vanish due to the boundary conditions. For impedance boundary conditions, we set $X\coloneqq H^{2}(\Omega)$ and define the sesquilinear form $a(\cdot,\cdot)$ to be
$a(u,v)\coloneqq$ $\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}-k^{2}(u,v)_{L^{2}(\Omega)}$ (13)
- $\alpha(\Delta u,\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)}+\alpha i\theta(\Delta u,v)_{L^{2}(\partial\Omega)}+\beta i\theta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},v)_{L^{2}(\partial\Omega)}$
- $\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\nabla v\cdot\bm{n})_{L^{2}(\partial\Omega)}-i\theta(u,v)_{L^{2}(\partial\Omega)},$
where several boundary terms have been substituted using the boundary conditions. An in-depth discussion arguing that the traces appearing in (12) and (13) are well-defined can be found in Appendix A.
3 Analysis of the nematic Helmholtz–Korteweg equation
We denote by $\mathcal{A}$ the operator induced by the sesquilinear forms defined in (11), (12), or (13), respectively. To show the well-posedness of (10), we show that the operator $\mathcal{A}$ is T-coercive in the sound soft case, and weakly T-coercive and injective in the sound hard and impedance cases. We split the operator into a T-coercive part $A\in\mathcal{L}(X)$ and a compact part $K\coloneqq\mathcal{A}-A$, where $K\in\mathcal{L}(X)$ varies depending on the boundary conditions. In particular, we define $A\in\mathcal{L}(X)$ as
$(Au,v)_{H^{2}(\Omega)}\coloneqq\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}-k(u,v)_{L^{2}(\Omega)}.$ (14)
Thus, $K\coloneqq\mathcal{A}-A$ only contains boundary terms, and $K=0$ in the case of sound soft boundary conditions.
To construct a $T$ operator that satisfies the desired properties, we consider the following eigenvalue problem: find $u\in H^{2}_{0}(\Omega)$, $\lambda\in\mathbb{C}$ such that
$e(u,v)\coloneqq\alpha(\Delta u,\Delta v)_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta v)_{L^{2}(\Omega)}+(\nabla u,\nabla v)_{L^{2}(\Omega)}=\lambda(u,v)_{L^{2}(\Omega)}$ (15)
for all $v\in H^{2}_{0}(\Omega)$. We show that the operator associated with (15) is self-adjoint.
Theorem 3.1.
The eigenvalue problem (15) is self-adjoint.
Proof.
Proving the self-adjointness of the operator associated with (15) is a matter of tedious integration by parts. We will omit most steps and only include salient ones. Let $u,v\in H^{2}_{0}(\Omega)$ be arbitrary. We begin observing that $\nabla\cdot\nabla[\mathbf{n}^{\top}(\mathbf{n}^{T}\nabla u)]$ can be expanded as
$n_{x}^{2}\partial_{x}^{4}u+n_{y}^{2}\partial_{y}^{4}u+(n_{x}+n_{y})^{2}\partial_{x}^{2}\partial_{y}^{2}u+2(n_{x}n_{y})\partial_{x}^{3}\partial_{y}u+2(n_{x}n_{y})\partial_{x}\partial_{y}^{3}u,$ (16)
where we assumed that $\mathbf{n}=(n_{x},n_{y})^{\top}$, since by assumption we know that $\nabla\mathbf{n}$ is identically zero. By integration by parts, we immediately observe that the first three terms in (16) are self-adjoint. Due to the regularity of the functions $u,v\in H^{2}_{0}(\Omega)$, we can swap the order of differentiation and integrate by parts to show that the last two terms in (16) are self-adjoint as well. The argument here presented can be easily extended to three dimensional case. ∎
Lemma 3.2.
For $\beta$ sufficiently small, the eigenvalue problem (15) is well-posed and the associated solution operator $S:L^{2}(\Omega)\to L^{2}(\Omega)$ is compact.
Proof.
Since $\|\mathbf{n}\|_{L^{\infty}(\Omega)}=1$, $\beta\geq 0$, and $\|\Delta v\|_{L^{2}(\Omega)}\leq\sqrt{d}|v|_{H^{2}(\Omega)}$, we have that
$|\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta u)_{L^{2}(\Omega)}|\leq\beta\|\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n}\|_{L^{2}(\Omega)}\|\Delta u\|_{L^{2}(\Omega)}\leq\beta\sqrt{d}|u|^{2}_{H^{2}(\Omega)}.$
Let $C_{P}>0$ be the constant from the Poincaré inequality, that is $\|u\|_{L^{2}(\Omega)}\leq C_{P}\|\nabla u\|_{L^{2}(\Omega)}$ for all $u\in H^{1}_{0}(\Omega)$. Then, for $\beta<\min\{\alpha,1/2,C_{P}^{2}/2\}$ we obtain that
$|e(u,u)|$ $=\alpha\|\Delta u\|^{2}_{L^{2}(\Omega)}+|\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\Delta u)_{L^{2}(\Omega)}|+\|\nabla u\|^{2}_{L^{2}(\Omega)}$ (17)
$\geq\alpha\|\Delta u\|^{2}_{L^{2}(\Omega)}-\beta\|u\|^{2}_{H^{2}(\Omega)}+\frac{1}{2}\|\nabla u\|^{2}_{L^{2}(\Omega)}+\frac{C_{P}^{2}}{2}\|u\|^{2}_{L^{2}(\Omega)}$
$\geq\left(\min\{\alpha,1/2,C_{P}^{2}/2\}-\beta\right)\|u\|^{2}_{H^{2}(\Omega)}\geq\gamma_{*}\|u\|^{2}_{H^{2}(\Omega)},$
and hence $e(\cdot,\cdot)$ is coercive, with $\gamma_{*}\coloneqq\min\{\alpha,1/2,C_{P}^{2}/2\}-\beta>0$. Here, we note that the Miranda–Talenti inequality [Grisvard, MPS00] gives that $|v|^{2}_{H^{2}(\Omega)}\leq\|\Delta v\|^{2}_{L^{2}(\Omega)}$ for all $v\in H^{2}_{0}(\Omega)$, since $\Omega$ is assumed to be convex. The compactness of the solution operator follows from the compact embedding $H^{2}_{0}(\Omega)\hookrightarrow L^{2}(\Omega)$. ∎
Remark 3.3 (Smallness assumption on $\beta$).
In applications, the parameters $\alpha$ and $\beta$ are usually small [VirgaNematoacoustics] and thus, the smallness assumption $\beta<\min\{\alpha/\sqrt{d},1/2,C_{P}^{2}/2\}$ essentially translates to $\beta<\alpha/\sqrt{d}$, because the Poincaré constant $C_{P}$ is generally larger than one. In fact, we usually have that $\beta\ll\alpha$ [VirgaNematoacoustics] so that the smallness assumption on $\beta$ is not restrictive for applications.
A direct consequence of Lemma 3.2 is that, for sufficiently small $\beta$, the eigenvalues of (15) are real, of finite multiplicity and that the eigenfunctions form an orthonormal basis of $L^{2}(\Omega)$ [Boffi2010]. We denote the eigenvalues by $\{\lambda^{(i)}\}_{i\in\mathbb{N}}$ and the corresponding eigenfunctions by $\{e^{(i)}\}_{i\in\mathbb{N}}$. Without loss of generality, we assume that the eigenfunctions are normalised with respect to the $H^{2}$-norm.
From this point on we will assume that $k^{2}\not\in\{\lambda^{(i)}\}_{i\in\mathbb{N}}$. Let $i_{\ast}$ be the largest index such that $\lambda^{(i)}<k^{2}$, i.e. $i_{\ast}\coloneqq\max\{i\in\mathbb{N}:\lambda^{(i)}<k^{2}\}$. Then, we define [^1] For readability, we neglect the trivial case where $k^{2}<\lambda^{(1)}$.
$W\coloneqq\text{span}_{0\leq i\leq i_{\ast}}\{e^{(i)}\},\qquad T\coloneqq\operatorname{id}_{X}-2P_{W}$ (18)
where $P_{W}:X\to W$ is the orthogonal projection onto $W$, orthogonal in the bilinear form on the left-hand side of (15). Since $T^{2}=\operatorname{id}$, the operator $T$ is bijective.
Theorem 3.4.
For $\beta$ sufficiently small, the operator $A\in\mathcal{L}(X)$ defined by (14) is T-coercive with respect to the $T$-operator defined in (18).
Proof.
Let $u\in X$ be arbitrary and set $u^{(i)}\in\mathbb{C}$ s.t. $u=\sum_{i\in\mathbb{N}}u^{(i)}e^{(i)}$. We note that $(Au,v)_{H^{2}(\Omega)}=e(u,v)-k^{2}(u,v)_{L^{2}(\Omega)}$ and calculate
$(ATu,u)_{H^{2}}$ $=\sum_{0\leq i\leq i_{\ast}}C(\lambda^{(i)})(k^{2}-\lambda^{(i)})(u^{(i)})^{2}+\sum_{i>i_{\ast}}C(\lambda^{(i)})(\lambda^{(i)}-k^{2})(u^{(i)})^{2}$
$\geq\gamma\|u\|^{2}_{X}$
with $\gamma\coloneqq\min_{i\in\mathbb{N}}\{C(\lambda^{(i)})|\lambda^{(i)}-k^{2}|\}>0$, where $C(\lambda^{(i)})$ is the renormalisation constant associated with the unitary condition of the eigenfunctions with respect to the $H^{2}$-norm. Thus, $A$ is indeed T-coercive. ∎
In the case of sound soft boundary conditions (11), this concludes the analysis of the continuous problem (10) as $K=0$. In the other cases, i.e. (12) and (13) we still have to show that the operator $K$ is compact and that the operator $\mathcal{A}$ is injective.
Lemma 3.5.
The operator $K\in\mathcal{L}(X)$ defined as $K=\mathcal{A}-A$, where $\mathcal{A}$ is the operator associated with (12) or (13), is compact.
Proof.
Let us consider first the case of sound hard boundary conditions. Let $\operatorname{tr}:X\to L^{2}(\partial\Omega)$ denote the trace operator. We recall that the boundary terms are well-defined and that we can use the compactness of the trace operator on Lipschitz domains, as discussed in Appendix A. We have that
$\|Ku\|_{H^{2}(\Omega)}=\sup_{v\in X\setminus\{0\}}\frac{|(Ku,v)_{H^{2}(\Omega)}|}{\|v\|_{H^{2}(\Omega)}}$ $\leq\sup_{v\in X\setminus\{0\}}\frac{|\alpha|\|\operatorname{tr}\Delta u\|_{L^{2}(\partial\Omega)}\|\operatorname{tr}\nabla v\cdot\bm{\nu}\|_{L^{2}(\partial\Omega)}}{\|v\|_{H^{2}(\Omega)}}$
$\leq C|\alpha|\|\operatorname{tr}\Delta u\|_{L^{2}(\partial\Omega)},$
where we used the continuity of the normal trace operator. Similarly, we obtain that
$|\beta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)}|$ $\leq\beta\|\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n}\|_{L^{2}(\partial\Omega)}\|\nabla v\cdot\bm{\nu}\|_{L^{2}(\partial\Omega)}$
$\leq C\beta\|\operatorname{tr}\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\|_{L^{2}(\partial\Omega)}\|\nabla v\|_{L^{2}(\Omega)}.$
For impedance boundary conditions, we calculate
$\|Ku\|_{H^{2}(\Omega)}$ $=\sup_{v\in X\setminus\{0\}}\frac{|(Ku,v)_{H^{2}(\Omega)}|}{\|v\|_{H^{2}(\Omega)}}$
$\leq\sup_{v\in X\setminus\{0\}}\frac{|\alpha|\|\operatorname{tr}\Delta u\|_{L^{2}(\partial\Omega)}(\|\operatorname{tr}\nabla v\cdot\bm{\nu}\|_{L^{2}(\partial\Omega)}+|\theta|\|\operatorname{tr}v\|_{L^{2}(\partial\Omega)})}{\lVert v\lVert_{H^{2}(\Omega)}}$
$+\sup_{v\in X\setminus\{0\}}\frac{|\theta|\|\operatorname{tr}u\|_{L^{2}(\partial\Omega)}\|\operatorname{tr}v\|_{L^{2}(\partial\Omega)}}{\|v\|_{H^{2}(\Omega)}}$
$\leq C(\|\operatorname{tr}\Delta u\|_{L^{2}(\partial\Omega)}+\|\operatorname{tr}u\|_{L^{2}(\partial\Omega)}),$
where we use the continuity of the normal trace operator and the compactness of the trace operator. Furthermore, we obtain for the nematic term that
$|$ $\beta i\theta(\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n},v)_{L^{2}(\partial\Omega)}-\beta(\mathbf{n}^{T}(\mathcal{H}u)\mathbf{n},\nabla v\cdot\bm{\nu})_{L^{2}(\partial\Omega)}|$
$\leq C\beta(1+|\theta|)\|\operatorname{tr}\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\|^{2}_{L^{2}(\partial\Omega)}\left(\|v\|_{L^{2}(\partial\Omega)}+\|\nabla v\|_{L^{2}(\Omega)}\right).$
Combining these estimates with the estimates from Lemma 3.5, we obtain for both cases that
$\|Ku\|_{H^{2}}\leq C\left(\|\operatorname{tr}\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\|_{L^{2}(\partial\Omega)}+\|\operatorname{tr}\Delta u\|_{L^{2}(\partial\Omega)}+\|\operatorname{tr}u\|_{L^{2}(\partial\Omega)}\right).$
Since the trace operator $\operatorname{tr}$ is compact, we obtain that $Ku_{n}\to Ku$ for any sequence $u_{n}\rightharpoonup u$ in $X$, and thus $K$ is compact. ∎
Lemma 3.6.
The operator $\mathcal{A}$ associated with (12) or (13) is injective.
Proof.
For sound hard boundary conditions this follows from the assumption that $k^{2}\not\in\{\lambda^{(i)}\}_{i\in\mathbb{N}}$. For the impedance case, we essentially adapt the argument from the Helmholtz case, cf. [EG_FE2, Thm. 35.5]. Taking $v\in\ker\mathcal{A}$ we obtain using Young’s inequality with $\zeta_{1},\zeta_{2}>0$
$0=|$- $\Im a(v,v)|=|-\alpha\theta(\Delta v,v)_{L^{2}(\partial\Omega)}-\beta\theta(\mathbf{n}^{\top}(\mathcal{H}v)\mathbf{n},\Delta v)_{L^{2}(\partial\Omega)}+\theta\|v\|_{L^{2}(\partial\Omega)}|$
$\geq\left|\frac{\alpha\zeta_{1}+\beta\theta\zeta_{2}}{2}\|\Delta v\|^{2}_{L^{2}(\partial\Omega)}+\frac{\beta\theta}{2\zeta_{2}}\|\mathbf{n}^{\top}(\mathcal{H}v)\mathbf{n}\|^{2}_{L^{2}(\partial\Omega)}+\frac{\theta}{2\zeta_{1}}\|v\|_{L^{2}(\partial\Omega)}^{2}\right|$
$\geq\left|\frac{\alpha\zeta_{1}+\beta\theta\zeta_{2}}{2}\|\Delta v\|^{2}_{L^{2}(\partial\Omega)}+\frac{\theta}{2\zeta}\|v\|_{L^{2}(\partial\Omega)}^{2}\right|.$
which implies that $\operatorname{tr}v=0$ and $\operatorname{tr}\Delta v=0$ on $\partial\Omega$. Let $B_{R}$ be a ball of radius $R$ such that $\Omega\subset B_{R}$ which we assume to be centered at zero and let $v^{\mathcal{E}}_{R}$ be the extension by zero of $v$ onto $B_{R}\cap\Omega^{c}$. Then, we have that $v^{\mathcal{E}}_{R}\in H^{2}_{0}(B_{R})$ and since $\operatorname{tr}\Delta v=0$ on $\partial\Omega$, we have that $\Delta v^{\mathcal{E}}_{R}=0$ in $B_{R}\cap\Omega^{c}$. Due to the Robin boundary condition, we have that $\nabla(\Delta v^{\mathcal{E}}_{R})\cdot\bm{\nu}|_{\partial\Omega}=0$. Hence $\nabla\Delta v$ is continuous across $\partial\Omega$. Since $(\nabla\Delta v^{\mathcal{E}}_{R})|_{\Omega}\in H(\operatorname{div},\Omega)$ and $(\nabla\Delta v^{\mathcal{E}}_{R})|_{B_{R}\cap\Omega^{c}}\in H(\operatorname{div},B_{R}\cap\Omega^{c})$, it follows that $\Delta^{2}v^{\mathcal{E}}_{R}\in L^{2}(B_{R})$. Since $v^{\mathcal{E}}_{R}$ is zero in an open subset of $B_{R}$, we can use the unique continuation principle for elliptic differential operators, cf. [HormanderPt3, Thm. 17.2.6], to conclude that $v^{\mathcal{E}}_{R}=0$ in $B_{R}$ and thus $v=0$ in $\Omega$. ∎
Altogether, we obtain that the operator $\mathcal{A}$ associated with (10) is weakly T-coercive and injective, and thus (10) is well-posed for all types of boundary conditions considered.
Remark 3.7 (Weak T-coercivity vs weak coercivity).
Instead of using weak T-coercivity, we could have shown that the operator $\mathcal{A}$ is weakly coercive (i.e. weakly T-coercive with $T=\operatorname{Id}$) by associating the term $k^{2}(u,v)_{L^{2}(\Omega)}$ to the compact perturbation. This approach would avoid the explicit construction of $T$ through the eigenvalue problem (15), but use the same arguments as in (17) to show coercivity. However, we chose to use weak T-coercivity because a discussion of the eigenvalue problem (15) is valuable in itself. For example, we later use the eigenvalue problem in Section 5.4 to study tunable resonators.
Remark 3.8 (Helmholtz–Korteweg equation).
Setting $\beta=0$ in (1), we obtain the Helmholtz–Korteweg equation. The continuous analysis presented in this section and the forthcoming discrete analysis allow for this case, yielding well-posedness of the Helmholtz–Korteweg equation and its discretisation. Furthermore, if we redefine $\|u\|_{X_{\alpha}}\coloneqq\alpha|u|_{H^{2}(\Omega)}+\|u\|_{H^{1}(\Omega)}$, we obtain continuity and stability of the sesquilinear form $a(\cdot,\cdot)$ with constants independent of $\alpha$. Thus, for $u_{0}\in H^{2}(\Omega)$ being a weak solution of the Helmholtz equation and $u_{\alpha}\in H^{2}(\Omega)$ being the solution of the Helmholtz–Korteweg equation with the same data, we obtain that
$\|u_{0}-u_{\alpha}\|_{X_{\alpha}}\leq\gamma^{-1}\sup_{w\in X}\frac{|a(u_{0}-u_{\alpha},w)|}{\|w\|_{X_{\alpha}}}\leq\alpha\gamma^{-1}\|\Delta u_{0}\|_{L^{2}(\Omega)}.$
Thus, taking $\alpha\to 0$ this yields that $u_{\alpha}\to u_{0}$. With similar arguments, we can obtain a similar result for the case $\beta\to 0$.
4 Discretization of the nematic Helmholtz–Korteweg equation
In this section, we analyse the discretisation of (10) using $H^{2}$-conforming finite elements. The analysis holds in two and three dimensions; our numerical experiments below will only be in two dimensions, as Firedrake [Firedrake] does not yet support such elements in three dimensions.
Conforming $H^{2}$ discretisations are naturally high-order, which is sometimes seen as a disadvantage; in this context it is appealing, since high-order discretisations achieve quasi-optimality for large wave numbers on much coarser meshes [Spence2025, DuWu2015, Babuska1995, Babuska1997]. For brevity we focus our analysis on the sound soft and the impedance boundary conditions; sound hard boundary conditions can be analysed similarly. Let $\{\mathcal{T}_{h}\}_{h}$ be a sequence of shape regular simplicial triangulations of $\Omega$. For polynomial degree $p\geq 3$, we define the finite element space $X_{h}\subset H^{2}(\Omega)$ as
$X_{h}\coloneqq\{v\in H^{2}(\Omega):v|_{T}\in\mathcal{P}^{p}(T)\quad\forall T\in\mathcal{T}_{h}\},$
where $\mathcal{P}^{p}(T)$ is the space of polynomials of degree $p$ on an element $T\in\mathcal{T}_{h}$. To implement the boundary conditions in the sound soft case, we use Nitsche’s method [Nit71]. To this end, we define the following terms:
$N^{\Delta^{2}}_{h}(u_{h},v_{h})$ $\coloneqq\alpha(\nabla(\Delta u_{h})\cdot\bm{\nu},v_{h})_{L^{2}(\partial\Omega)}+\alpha(u_{h},\nabla(\Delta v_{h})\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$ (19a)
$N_{h}^{\beta}(u_{h},v_{h})$ $\coloneqq\beta(\nabla(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n})\cdot\bm{\nu},v_{h})_{L^{2}(\partial\Omega)}+\beta(u_{h},\nabla(\mathbf{n}^{T}(\mathcal{H}v_{h})\mathbf{n})\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$ (19b)
$N^{\Delta}_{h}(u_{h},v_{h})$ $\coloneqq(\nabla u_{h}\cdot\bm{\nu},v_{h})_{L^{2}(\partial\Omega)}+(u_{h},\nabla v_{h}\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$ (19c)
$S_{h}(u_{h},v_{h})$ $\coloneqq\alpha\eta_{1}h^{-3}(u_{h},v_{h})_{L^{2}(\partial\Omega)}+\beta\eta_{2}h^{-3}(u_{h},v_{h})_{L^{2}(\partial\Omega)}+\eta_{3}h^{-1}(u_{h},v_{h})_{L^{2}(\partial\Omega)},$ (19d)
where $\eta_{1},\eta_{2},\eta_{3}>0$ are suitably chosen stabilization parameters. Then, we consider the discrete problem: find $u_{h}\in X_{h}$ such that
$a_{h}(u_{h},v_{h})=(f,v_{h})_{L^{2}(\Omega)}\quad\forall v_{h}\in X_{h},$ (20)
where the sesquilinear form $a_{h}(\cdot,\cdot)$ is defined through
$a_{h}(u_{h},v_{h})\coloneqq a(u_{h},v_{h})+\epsilon\big($ $N^{\Delta^{2}}_{h}(u_{h},v_{h})-N^{\Delta}_{h}(u_{h},v_{h})+N^{\beta}_{h}(u_{h},v_{h})+S_{h}(u_{h},v_{h})\big).$ (21)
Here $a(\cdot,\cdot)$ is the sesquilinear form defined in (11) or (13) and $\epsilon=1$ in the sound soft case and $\epsilon=0$ in the impedance case. Due to the Nitsche terms, $a_{h}(\cdot,\cdot)\not=a(\cdot,\cdot)$ and $X_{h}\not\subset X$ if $\epsilon=1$, so our discretisation is non-conforming in the sound soft case. To ensure that the sesquilinear form $a_{h}(\cdot,\cdot)$ is well-defined for elements $u\in X$, we require more regularity since we have to evaluate the Nitsche terms defined in (19a) and (19b). Thus, we make the following assumption.
Assumption 4.1.
The solution $u$ of the continuous problem (10) with sound soft boundary conditions satisfies $u\in\tilde{X}\coloneqq X\cap H^{5/2}(\Omega)$.
To show stability of (20), we follow the strategy from the continuous case. Thus, we consider the discretisation of the eigenvalue problem (15): find $u_{h}\in\tilde{X}_{h}\subseteq X_{h}$, $\lambda\in\mathbb{R}$ s.t.
$e_{h}(u_{h},v_{h})\coloneqq$ $\alpha(\Delta u_{h},\Delta v_{h})_{L^{2}(\Omega)}+\beta(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n},\Delta v_{h})_{L^{2}(\Omega)}+(\nabla u_{h},\nabla v_{h})_{L^{2}(\Omega)}$ (22)
$+$ $\epsilon\left(N^{\Delta^{2}}_{h}(u_{h},v_{h})+N^{\beta}_{h}(u_{h},v_{h})-N^{\Delta}_{h}(u_{h},v_{h})+S_{h}(u_{h},v_{h})\right)=\lambda(u_{h},v_{h})_{L^{2}(\Omega)}$
for all $v_{h}\in\tilde{X}_{h}\subseteq X_{h}$. For the impedance case we assume that the boundary conditions $u_{h}=0$ and $\Delta u_{h}=0$ on $\partial\Omega$ are imposed through the discrete space $\tilde{X}_{h}$ and naturally, respectively. For the sound soft case, we have that $\tilde{X}_{h}=X_{h}$ and the boundary conditions are imposed through the Nitsche terms (19).
Remark 4.2 (On Nitsche’s method).
The main motivation for using Nitsche’s method is the difficulty of imposing essential boundary conditions for $C^{1}$-conforming finite elements, cf. [Kirby]. We note that the forthcoming analysis includes the case where this is achieved by setting $\epsilon=0$ and modifying $\tilde{X}_{h}$ also in the sound soft case.
4.1 Analysis of the discrete eigenvalue problem
Before considering the approximation of (10), we analyse the discrete eigenvalue problem (22). In preparation for showing uniform coercivity of $e_{h}(\cdot,\cdot)$, we require the following inverse trace inequality.
Lemma 4.3.
There exists a constant $C>0$ such that the following inverse trace inequality holds
$\|\nabla(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n})\cdot\bm{\nu}\|^{2}_{L^{2}(\partial\Omega)}\leq Ch^{-3}|u_{h}|^{2}_{H^{2}(\Omega)}\qquad\forall u_{h}\in\tilde{X}_{h}.$ (23)
Proof.
We set $\mathcal{K}_{h}\coloneqq\{v_{h}\in tilde{X}_{h}:\nabla(\mathbf{n}^{\top}\mathcal{H}v_{h}\mathbf{n})=0\}$ such that $\tilde{X}_{h}=\mathcal{K}_{h}^{\perp}\oplus\mathcal{K}_{h}$ and consider the following eigenvalue problem on $\mathcal{K}_{h}^{\perp}$: find $u_{h}\in\mathcal{K}_{h}^{\perp}$, $\tilde{\lambda}_{h}\in\mathbb{R}$ such that for all $v_{h}\in\mathcal{K}_{h}^{\perp}$ we have
$h(\nabla(\mathbf{n}^{\top}\mathcal{H}u_{h}\mathbf{n})\cdot\bm{\nu},\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n})\cdot\bm{\nu})_{L^{2}(\partial\Omega)}=\tilde{\lambda}_{h}(\nabla(\mathbf{n}^{T}\mathcal{H}u_{h}\mathbf{n}),\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n}))_{L^{2}(\Omega)}.$ (24)
The left hand-side is bounded and the right hand-side is coercive on $\mathcal{K}_{h}^{\perp}$, so the problem is well-posed, and the associated eigenvalues are positive and finite. Further, the min-max characterisation yields for the maximal eigenvalue $\tilde{\lambda}_{h,\max}$ that
$\tilde{\lambda}_{h,\max}=\sup_{v_{h}\in\mathcal{K}_{h}^{\perp}\setminus\{0\}}\frac{h(\nabla(\mathbf{n}^{\top}\mathcal{H}v_{h}\mathbf{n})\cdot\bm{\nu},\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n})\cdot\bm{\nu})_{L^{2}(\partial\Omega)}}{(\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n}),\nabla(\mathbf{n}^{T}\mathcal{H}v_{h}\mathbf{n}))_{L^{2}(\Omega)}}.$ (25)
Thus, we obtain that $\|\nabla(\mathbf{n}^{\top}\mathcal{H}u_{h}\mathbf{n})\cdot\bm{\nu}\|_{L^{2}(\partial\Omega)}^{2}\leq Ch^{-1}\|\nabla(\mathbf{n}^{T}\mathcal{H}u_{h}\mathbf{n}^{T})\|_{L^{2}(\Omega)}^{2}$. Then, we use the a standard inverse inequality and that $\|\mathbf{n}\|^{2}_{L^{\infty}}=1$ to estimate
$h^{-1}\|\nabla(\mathbf{n}^{\top}\mathcal{H}u_{h}\mathbf{n}^{T})\|_{L^{2}(\Omega)}^{2}\leq Ch^{-3}\|\mathcal{H}u_{h}\|_{L^{2}(\Omega)}^{2}\leq Ch^{-3}|u_{h}|_{H^{2}(\Omega)}^{2}.$
Lemma 4.4.
There exist constants $C^{\Delta^{2}},C^{\beta},C^{\Delta}>0$ s.t. for all $u_{h}\in\tilde{X}_{h}$ and all $\zeta_{1},\zeta_{2},\zeta_{3}>0$, we have the following estimates
[…]
$N_{h}^{\Delta}(u_{h},u_{h})$ $\leq C^{\Delta}\zeta_{3}h^{-1}\|\nabla u_{h}\|^{2}_{L^{2}(\Omega)}+\zeta_{3}^{-1}\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}.$
Proof.
First of all, we note that the following inequality holds [BET24, Lem. 5.2]:
$\|\nabla(\Delta u_{h})\cdot\bm{\nu}\|^{2}_{L^{2}(\partial\Omega)}\leq Ch^{-3}\|\Delta u_{h}\|^{2}_{L^{2}(\Omega)}.$
Combining this estimate with Young’s inequality, we obtain for all $\zeta_{1}>0$
$\alpha^{-1}N_{h}^{\Delta^{2}}(u_{h},u_{h})=2(\nabla(\Delta u_{h})\cdot\bm{\nu},u_{h})_{L^{2}(\partial\Omega)}\leq C\zeta_{1}h^{-3}\|\Delta u_{h}\|^{2}_{L^{2}(\Omega)}+\zeta_{2}^{-1}\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}.$
The second inequality follows with the same argument and the inverse trace inequality from Lemma 4.3, whereas the third inequality requires the standard inverse inequality. ∎
Let us introduce the following norm on the discrete space $X_{h}$:
$\|u_{h}\|_{\epsilon}^{2}\coloneqq|u_{h}|^{2}_{H^{2}(\Omega)}+\lvert u_{h}\rvert^{2}_{H^{1}(\Omega)}+\epsilon\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}.$ (26)
Notice that in the case of impedance boundary conditions, i.e. $\epsilon=0$, we consider the eigenvalue problem (22) to be posed on the space $\tilde{X}_{h}\cap H^{1}_{0}(\Omega)$, and consequently we can use the Poincaré inequality to show that the expression (26) is indeed a norm.
Lemma 4.5.
For $\beta$ sufficiently small, and $\eta_{1},\eta_{2},\eta_{3}>0$ sufficiently large, the bilinear form $e_{h}(\cdot,\cdot)$ is uniformly coercive on $X_{h}$, i.e. there exists a constant $\gamma>0$ independent of $h$ s.t.
$\lvert e_{h}(u_{h},u_{h})\rvert\geq\gamma\|u_{h}\|_{\epsilon}^{2}\quad\forall u_{h}\in\tilde{X}_{h}.$ (27)
Proof.
For $u_{h}\in X_{h}$ we estimate with the results from Lemma 4.4:
$\left|\epsilon\left(N_{h}^{\Delta^{2}}(u_{h},u_{h})+N_{h}^{\beta}(u_{h},u_{h})-N_{h}^{\Delta}(u_{h},u_{h})\right)\right|$
$\geq$ $\epsilon\Big(-C^{\Delta^{2}}\zeta_{1}\alpha h^{-3}\|\Delta u_{h}\|_{L^{2}(\Omega)}^{2}-C^{\beta}\zeta_{2}\beta h^{-3}|u_{h}|_{H^{2}(\Omega)}-C^{\Delta}\zeta_{3}h^{-1}\|\nabla u_{h}\|^{2}_{L^{2}(\Omega)}$
[…]
In the proof of Lemma 3.2, we used the fact that $|v|_{H^{2}(\Omega)}\leq\|\Delta v\|_{L^{2}(\Omega)}$ for all $v\in H^{2}_{0}(\Omega)$. In the impedance case ($\epsilon=0$), we can use that estimate since $\tilde{X}_{h}\subset H^{2}_{0}(\Omega)$, but in the sound soft case ($\epsilon=1$) no boundary conditions are imposed through the space $\tilde{X}_{h}\subset H^{2}(\Omega)$. However, elliptic regularity [Grisvard, Sec. 2.3.3] grants that $|v|_{H^{2}(\Omega)}\leq C(\|\Delta v\|_{L^{2}(\Omega)}+\|v\|_{H^{1}(\Omega)})$ for all $v\in H^{2}(\Omega)$ with constant $C>0$. Due to the compactness of the embedding $H^{2}(\Omega)\hookrightarrow H^{1}(\Omega)$, we may interchange the $H^{1}$-term with any term that guarantees injectivity on $H^{2}(\Omega)$ due to the Peetre–Tartar theorem [EG_FE1, Lem. A.20]. Thus, we obtain altogether that
$|v|_{H^{2}(\Omega)}^{2}\leq(1+\epsilon C_{\mathcal{H}})\|\Delta v\|_{L^{2}(\Omega)}^{2}+\epsilon C_{\mathcal{H}}\|v\|_{L^{2}(\partial\Omega)}^{2}\qquad\forall v\in H^{2}(\Omega),$
with constant $C_{\mathcal{H}}>0$.
Consequently, we obtain with $|(\beta\mathbf{n}^{\top}\mathcal{H}u_{h}\mathbf{n},\Delta u_{h})_{L^{2}(\Omega)}|\leq\beta\sqrt{d}|u_{h}|^{2}_{H^{2}(\Omega)}$ that
$|e_{h}(u_{h},u_{h})|\geq$ $\alpha\left(1-\epsilon C^{\Delta^{2}}\zeta_{1}h^{-3}\right)\|\Delta u_{h}\|^{2}_{L^{2}(\Omega)}-\beta\left(\sqrt{d}+\epsilon C^{\beta}\zeta_{2}h^{-3}\right)|u_{h}|^{2}_{H^{2}(\Omega)}$
$+\left(1-\epsilon C^{\Delta}\zeta_{2}h^{-1}\right)\|\nabla u_{h}\|^{2}_{L^{2}(\Omega)}+\epsilon\alpha\left(\eta_{1}h^{-3}-\zeta_{1}^{-1}\right)\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}$
$+\epsilon\beta\left(\eta_{2}h^{-3}-\zeta_{2}^{-1}\right)\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}+\epsilon\left(\eta_{3}h^{-1}-\zeta_{3}^{-1}\right)\|u_{h}\|^{2}_{L^{2}(\partial\Omega)}$
$\geq$ $\left(\alpha\left(1-\epsilon C^{\Delta^{2}}\zeta_{1}h^{-3}\right)(1+\epsilon C_{\mathcal{H}})^{-1}-\beta\left(\sqrt{d}+\epsilon C^{\beta}\zeta_{2}h^{-3}\right)\right)|u_{h}|_{H^{2}(\Omega)}$
$+\left(1-\epsilon C^{\Delta}\zeta_{2}h^{-1}\right)\|\nabla u_{h}\|^{2}_{L^{2}(\Omega)}+\epsilon\Big(\alpha\left(\eta_{1}h^{-3}-\zeta_{1}^{-1}-\left(1-\epsilon C^{\Delta^{2}}\zeta_{1}h^{-3}\right)\frac{\epsilon C_{\mathcal{H}}}{1+\epsilon C_{\mathcal{H}}}\right)$
$+\beta\left(\eta_{2}h^{-3}-\zeta_{2}^{-1}\right)+\left(\eta_{3}h^{-1}-\zeta_{3}^{-1}\right)\Big)|u_{h}|^{2}_{L^{2}(\partial\Omega)}$
$\geq$ $\,\gamma\|u_{h}\|^{2}_{\epsilon},$
where we assume that $\beta$ is sufficiently small, choose $\zeta_{1},\zeta_{2},\zeta_{3}$ sufficiently small and assume that the penalty parameters $\eta_{1},\eta_{2},\eta_{3}$ are chosen sufficiently large such that $\gamma>0$ independent of h. ∎
Remark 4.6 (Smallness assumption on $\beta$).
In the proof of Lemma 4.5, we require that
$\beta<\left(\alpha\left(1-\epsilon C^{\Delta^{2}}\zeta_{1}h^{-3}\right)(1+\epsilon C_{\mathcal{H}})^{-1}\right)\left(\sqrt{d}+\epsilon C^{\beta}\zeta_{2}h^{-3}\right)^{-1}.$
In the impedance case, where $\epsilon=0$, this translates to $\beta<\alpha/\sqrt{d}$. For the sound soft case, we note that $\zeta_{1},\zeta_{2}$ may be chosen arbitrarily small such that we essentially obtain the smallness assumption
$\beta<\alpha/(\sqrt{d}(1+\epsilon C_{\mathcal{H}})),$ (28)
but then penalty parameters $\eta_{1},\eta_{2}$ have to be chosen arbitrarily large. To obtain stability without having to choose the penalty parameters sufficiently large, we could modify the Nitsche formulation in the spirit of [L16], where the Nitsche terms are represented by suitable lifting operators. Then, we would obtain (28), but the modification would introduce more technicalities. For ease of presentation, we focus on the classical Nitsche formulation.
4.2 Analysis of the discrete problem
In addition to the norm $\|\cdot\|_{\epsilon}$ in (26), we define the stronger norm
[…]
Due to the inverse inequalities applied in Lemma 4.4, we have that $\|u_{h}\|^{2}_{\epsilon}\simeq\|u_{h}\|^{2}_{h,\epsilon}$ for all $u_{h}\in X_{h}$ with constants independent of $h$. The norm $\|\cdot\|_{h,\epsilon}$ is well-defined for elements $u\in\tilde{X}$, and there exists a constant $C_{\text{cont}}>0$ such that
$a_{h}(w,u_{h})\leq C_{\text{cont}}\|w\|_{h,\epsilon}\|v_{h}\|_{h,\epsilon}\quad\forall w\in\tilde{X}+X_{h},v_{h}\in X_{h}.$ (30)
In the following, we denote by $\mathcal{A}_{h}\in\mathcal{L}(X_{h})$ the operator associated with the sesquilinear form $a_{h}(\cdot,\cdot)$. For the impedance case, we define $K_{h}\in\mathcal{L}(X_{h})$ as
$(K_{h}u_{h},v_{h})_{H^{2}}\coloneqq$ $-\alpha(\Delta u_{h},\nabla v_{h}\cdot\bm{\nu})_{L^{2}(\partial\Omega)}+\alpha i\theta(\Delta u_{h},v_{h})_{L^{2}(\partial\Omega)}-i\theta(u_{h},v_{h})_{L^{2}(\partial\Omega)}$
$+\beta i\theta(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n},v_{h})_{L^{2}(\partial\Omega)}-\beta(\mathbf{n}^{\top}(\mathcal{H}u_{h})\mathbf{n},\nabla v_{h}\cdot\bm{\nu})_{L^{2}(\partial\Omega)},$
and set $A_{h}\coloneqq\mathcal{A}_{h}-K_{h}\in\mathcal{L}(X_{h})$. Note that $K_{h}\coloneqq\Pi_{h}K|_{X_{h}}$, where $\Pi_{h}:X\to X_{h}$ is the orthogonal projection onto $X_{h}$. For the sound soft case, we set $K_{h}=0$ such that $A_{h}\coloneqq\mathcal{A}_{h}\in\mathcal{L}(X_{h})$.
To show the well-posedness of the discrete problem (20), we will show that $\mathcal{A}_{h}$ is uniformly Th-coercive on $X_{h}$. In particular, it suffices to show that $A_{h}=\mathcal{A}_{h}-K$ is uniformly Th-coercive on $X_{h}$. In the impedance case, this statement is nontrivial and originates from the fact that the compact perturbation $K$ can be neglected asymptotically. The following theorem [^2] Note that this theorem implicitly uses the fact that the sequence of discrete spaces $(X_{h})_{h>0}$ is dense in $X$, i.e. that the approximation property is satisfied. formalises this intuition and since it is only applied in the impedance case, we state it in a conforming setting. We refer to [HLS22, Thm. 3] for a more general result and to [SBH19, Sec. 8.9] or [BHP17] for similar discussion on compactly perturbed coercive problems.
Theorem 4.7.
Let $\mathcal{A}\in\mathcal{L}(X)$ be injective and weakly T-coercive, that is $\mathcal{A}=A+K$, where $A\in\mathcal{L}(X)$ is T-coercive and $K\in\mathcal{L}(X)$ is compact. If there exists a family of bijective operators $(T_{h})_{h>0}$, $T_{h}\in\mathcal{L}(X_{h})$, such that $A$ is uniformly $T_{h}$-coercive on $X_{h}$, then there exists $h_{0}>0$ and a family of bijective operators $(\tilde{T}_{h})_{h\leq h_{0}}$, $\tilde{T}_{h}\in\mathcal{L}(X_{h})$, such that $\mathcal{A}$ is uniformly $\tilde{T}_{h}$-coercive on $X_{h}$ for all $h\leq h_{0}$.
Proof.
See e.g. [EG_FE2, Ex. 26.5] which is based on [W86]. For completeness, we provide a proof in Appendix B. ∎
To define a discrete analogue of $T$, let $\{\lambda^{(i)}_{h}\}$ and $\{e^{(i)}_{h}\}$ be the discrete eigenvalues and eigenfunctions obtained from the eigenvalue problem (22). As in the continuous case, we renormalise the eigenfunctions with respect to the $\|\cdot\|_{\epsilon}$-norm. We define the discrete operator $T_{h}\in\mathcal{L}(X_{h})$ as
$W_{h}\coloneqq\text{span}_{0\leq i\leq i_{\ast}}\{e^{(i)}_{h}\},\qquad T_{h}\coloneqq\operatorname{id}_{X_{h}}-2P_{W_{h}},$
where $P_{W_{h}}:X_{h}\to W_{h}$ is the orthogonal projection onto $W_{h}$. As before, we have that $T_{h}^{2}=\operatorname{id}$, thus $T_{h}$ is bijective. Notice that we are assuming that $h$ is small enough such that $\lambda^{(i_{\ast})}_{h}<k^{2}$.
Theorem 4.8.
For $h$ small enough such that $\lambda^{(i_{\ast})}_{h}<k^{2}$, $\beta$ sufficiently small and $\eta_{1},\eta_{2},\eta_{3}>0$ sufficiently large, the operator $A_{h}\in\mathcal{L}(X_{h})$ is uniformly Th-coercive on $X_{h}$.
Proof.
For any $u_{h}\in X_{h}$, we have that due to the smallness assumption on $h$
$(A_{h}T_{h}u_{h},u_{h})_{H^{2}}$ $=\sum_{0\leq i\leq i_{\ast}}C(\lambda^{(i)}_{h})(k^{2}-\lambda^{(i)}_{h})(u_{h}^{(i)})^{2}+\sum_{i>i_{\ast}}C(\lambda^{(i)}_{h})(\lambda^{(i)}_{h}-k^{2})(u_{h}^{(i)})^{2}$
$\geq C(\epsilon)\gamma\|u_{h}\|^{2}_{\epsilon},$
where $\gamma\coloneqq\min_{i\in\mathbb{N}}C(\lambda^{(i)}_{h})\{|\lambda^{(i)}_{h}-k^{2}|\}>0$ is uniform in $h$, with $C(\lambda^{(i)}_{h})$ being the renormalisation constant. Thus $A_{h}$ is uniformly Th-coercive on $X_{h}$. ∎
For sound soft boundary conditions, this immediately shows that the discrete problem is well-posed, since $\mathcal{A}_{h}=A_{h}$. For impedance boundary conditions, we can apply Theorem 4.7 to infer the asymptotic stability of $\mathcal{A}_{h}$ from the stability of $A_{h}$.
Corollary 4.9.
Let $u\in\tilde{X}$ be the solution to (10) and $u_{h}\in X_{h}$ be the solution to the discrete problem (20). Then, it holds that
$a_{h}(u-u_{h},v_{h})=0\quad\forall v_{h}\in X_{h}.$
Proof.
For $\epsilon=0$, the statement is trivial since $a_{h}(\cdot,\cdot)=a(\cdot,\cdot)$. In the sound soft case ($\epsilon=1$), the statement follows from partial integration as usual for the Nitsche’s method, where we recall that $u\in\tilde{X}$ has enough regularity to ensure that the Nitsche terms (19) are well-defined. ∎
Thus, we have the following best approximation result.
Theorem 4.10.
Let $u\in\tilde{X}$ be the solution to (10). Assuming that $\eta_{1},\eta_{2},\eta_{3}>0$ are sufficiently large, there exists $h_{0}>0$ such that the discrete problem (20) has a unique solution $u_{h}\in X_{h}$ for all $h<h_{0}$. Further, there exists a constant $C>0$ such that
$\|u-u_{h}\|_{\epsilon}\leq C\inf_{v_{h}\in X_{h}}\|u-v_{h}\|_{h,\epsilon}.$
Proof.
Due to Theorem 4.8, the operator $A_{h}\in L(X_{h})$ is uniformly $T_{h}$-coercive on $X_{h}$ for $h<h_{\ast}$, where $h_{\ast}\coloneqq\max_{h}\{\lambda_{h}^{(i_{\ast})}<k^{2}\}$. Due to Theorem 4.7, there further exists $\tilde{h}_{\ast}\leq h_{\ast}$ such that $\mathcal{A}_{h}=A_{h}+K_{h}$ is uniformly $\tilde{T}_{h}$-coercive on $X_{h}$ for $h<\tilde{h}_{\ast}$. Thus, the discrete problem (20) has a unique solution $u_{h}\in X_{h}$ for all $h<h_{0}$, where $h_{0}=h_{\ast}$ if $\epsilon=1$ and $h_{0}=\tilde{h}_{\ast}$ if $\epsilon=0$.
For $v_{h}\in X_{h}$ the triangle inequality yields that
$\|u-u_{h}\|_{\epsilon}\leq\|u-v_{h}\|_{\epsilon}+\|v_{h}-u_{h}\|_{\epsilon}.$
Using the uniform $T_{h}$-coercivity (where we rename $\tilde{T}_{h}$ as $T_{h}$ in the case of impedance boundary conditions) of $a_{h}(\cdot,\cdot)$ and Corollary 4.9, we obtain for the second term,
$\gamma\|u_{h}-v_{h}\|_{\epsilon}^{2}$ $\leq a_{h}(T_{h}(u_{h}-v_{h}),u_{h}-v_{h})\leq a_{h}(u-v_{h},T_{h}^{\ast}(u_{h}-v_{h}))$
$\leq C_{\text{cont}}\|T_{h}\|_{\mathcal{L}(X_{h})}\|u-v_{h}\|_{h,\epsilon}\|u_{h}-v_{h}\|_{h,\epsilon}.$
We note that $\|v_{h}\|_{\epsilon}\leq\|v_{h}\|_{h,\epsilon}$, divide by $\|u_{h}-v_{h}\|_{h,\epsilon}$, and take the infimum over $v_{h}\in X_{h}$ we obtain
$\|u-u_{h}\|_{\epsilon}\leq\left(1+\frac{C_{\text{cont}}\|T_{h}\|_{\mathcal{L}(X_{h})}}{\gamma}\right)\inf_{v_{h}\in X_{h}}\|u-v_{h}\|_{h,\epsilon}.$
Remark 4.11 (The threshold $h_{0}$).
The threshold $h_{0}$ is Theorem 4.10 is in general not explicitly given. However, for sound soft boundary conditions, it can be characterised by the condition $\lambda^{(i_{\ast})}_{h}<k^{2}$. We could adapt a similar scheme as considered in [vBZ24] to ensure that the smallness assumption on $h$ is satisfied: on a sequence of refined meshes, the criterion $\lambda^{(i_{\ast})}_{h}<k^{2}$ is checked by solving the eigenvalue problem (22). If the criterion is met, the assumptions of Theorem 4.8 are satisfied, otherwise the mesh is refined and the procedure is repeated. To reduce the required number of degrees of freedom, an adaptive error estimator for the first $i_{\ast}$ eigenfunctions can be used. In the case of impedance boundary conditions, this argument is not applicable.
Returning to Remark 3.7, we note that a similar characterisation of $h_{0}$ can be obtained in the sound soft case using a weak coercivity argument by employing the techniques described in [D94]. Therein, $h_{0}$ is characterised by the sufficient approximation of the eigenvalues of the compact term $A^{-1}K$ by those of $A_{h}^{-1}K_{h}$. In our setting, this leads to the same condition $\lambda^{(i_{\ast})}_{h}<k^{2}$. It might be possible to transfer the arguments from [D94] from the weakly coercive to the weakly T-coercive case by studying the approximation of $(AT)^{-1}K$ by $(A_{h}T_{h})^{-1}K_{h}$. This might allow the study of the impedance case as well.
Remark 4.12 (Sound hard boundary conditions).
The case of sound hard boundary conditions can be analysed similarly. To implement the boundary conditions with Nitsche’s method, we set $N^{\Delta^{2}}_{h}=0$ and redefine the penalization term as
$S_{h}(u_{h},v_{h})\coloneqq\frac{\eta}{h}(\nabla u_{h}\cdot\bm{\nu},\nabla v_{h}\cdot\bm{\nu})_{L^{2}(\partial\Omega)}.$
The analysis requires the study of a different discrete and continuous eigenvalue problem with respect to (15) and (22). This analysis is similar and we omit the details.
5 Numerical examples
The analysis above is abstract and applies to any $C^{1}$ conforming finite element space. In two dimensions these include the Argyris element [Argyris] ($p\geq 5$), the Hsieh–Clough–Tocher macroelement defined on the Alfeld split of a triangular mesh [HCT] ($p\geq 3$), and the Morgan–Scott element [MorganScott, Parker] ($p=5$). Here we will present numerical simulations implemented using the Firedrake finite element library and ngsPETSc [Firedrake, brubeck2025a, ngsPETSc]. We test the Argyris element and the Hsieh–Clough–Tocher macroelements.
5.1 Quasi optimal convergence
Plane waves are solutions of the Helmholtz equation that remain constant over a plane perpendicular to the direction of propagation. They are a natural choice for the numerical validation of the nematic Helmholtz–Korteweg equation, since they are also solutions of the latter, i.e.
$u(\mathbf{x})=\exp\left(i(\mathbf{d}\cdot\mathbf{x})\right),\qquad\mathbf{d}\in\mathbb{C}.$ (31)
Substituting (31) into (1), it is easy to observe that plane waves are also exact solution of the nematic Helmholtz–Korteweg equation, under appropriate choice of the wave vector $\mathbf{d}$ [farrell2024b]. In particular, if $d\coloneqq|\mathbf{d}|$ satisfies the following dispersion relation (32), then the plane wave is also a solution of the nematic Helmholtz–Korteweg equation:
$\alpha d^{4}+\beta d^{2}(\mathbf{d}^{\top}\mathbf{n})^{2}+d^{2}-k^{2}=0.$ (32)
In Figure 1 we present the convergence of the $H^{2}$-norm of the error for the nematic Helmholtz–Korteweg equation with boundary data constructed in such a way that the exact solution is a plane wave. We observe that from Theorem 4.10 the error should converge with a rate only determined by the approximation properties of the finite element space. In particular, it is well known, see for example [BrennerScott, Section 5.9], that one can construct an interpolation operator from $H^{5}(\Omega)$ to the Argyris finite element space that converges with rate $h^{3}$ with respect to the $H^{2}$-norm [Argyris], i.e.
$\|u-I_{h}u\|_{H^{2}(\Omega)}\leq Ch^{4}\|u\|_{H^{5}(\Omega)}.$ (33)
Thus, combining (33) with Theorem 4.10 we expect the error to converge with rate $h^{4}$ with respect to the $H^{2}$-norm, as confirmed by Figure 1. Similarly, if we consider the Hsieh–Clough–Tocher macroelement [HCT] we expect the error to converge with rate $h^{2}$ with respect to the $H^{2}$-norm, as confirmed by Figure 1.
[…]
Figure 1: The convergence of the $H^{2}$-norm of the error for the nematic Helmholtz–Korteweg equation for different values of $k$ (top row) and the corresponding manufactured solution (bottom row),where we display the real part of the condensation wave $u(\mathbf{x})$, i.e. $p(\mathbf{x},t)=\rho_{0}\left(1+\Re\{u(\mathbf{x})e^{-i\omega t}\}\right)$, with $\rho(\mathbf{x},t)$ being the density at position $\mathbf{x}$ and time $t$ and $\rho_{0}$ the mean density of the fluid at rest.
5.2 Anisotropic Gaussian pulse
We study the anisotropic effect of the nematic term $\beta\nabla\cdot\nabla\left(\mathbf{n}^{\top}\mathcal{H}u\mathbf{n}\right)$ on the propagation of a symmetric Gaussian pulse of the form
$f(x,y)=\exp\left(-(40^{2})\left[\left(x-\frac{1}{2}\right)^{2}+\left(y-\frac{1}{2}\right)^{2}\right]\right).$ (34)
Several authors have studied the propagation of plane waves in the context of nematic liquid crystals [farrell2024b, VirgaNematoacoustics, SelingerEtAll]. In particular, it has been shown that the nematic term causes an anisotropic speed of propagation of plane waves, which is greater in the direction of the nematic director $\mathbf{n}$. In Figure 2 we present numerical simulations of the propagation of a symmetric Gaussian pulse by the nematic Helmholtz–Korteweg equation with sound soft boundary conditions, and in Figure 3 we impose impedance boundary conditions. Both figures show that the anisotropic speed of propagation observed for plane waves is also present for the Gaussian pulse, independent of the choice of boundary conditions.
((a))
((b))
((c))
((d))
Figure 2: The propagation of a symmetric Gaussian pulse by the nematic Helmholtz–Korteweg equation with sound soft boundary conditions when $\beta=0$ (a) and when $\beta=1\cdot 10^{-2}$ and $\mathbf{n}$ is aligned with the $x$-axis (b), the diagonal (c) and the $y$-axis (d). The parameters here are $k=40$, $\alpha=10^{-2}$.
((a))
((b))
((c))
((d))
Figure 3: The propagation of a symmetric Gaussian pulse by the nematic Helmholtz–Korteweg equation with impedance boundary conditions when $\beta=0$ (a) and when $\beta=1\cdot 10^{-2}$ and $\mathbf{n}$ is aligned with the $x$-axis (b), the diagonal (c) and the $y$-axis (d). The parameters here are $k=40$, $\alpha=10^{-2}$.
5.3 The Mullen–Lüthi–Stephen experiment
The Mullen–Lüthi–Stephen experiment [MullenLuti] consists of a planar acoustic wave propagating in a nematic liquid crystal, where the nematic director $\mathbf{n}$ kept fixed in the $x$-direction via a magnetic field. The experiment reveals that the speed of propagation of the acoustic wave is anisotropic, being greater in the direction of the nematic director $\mathbf{n}$.
We consider a variant of the Mullen–Lüthi–Stephen experiment, where the nematic director $\mathbf{n}$ is kept fixed in the $x$-direction in the central region of the domain and in the $y$-direction in the outer region of the domain. We then study the propagation of a planar acoustic wave coming from the top of the domain. We can clearly see from Figure 4 that the speed of propagation of the acoustic wave is anisotropic, being greater in the central region of the domain where the nematic director $\mathbf{n}$ is aligned with the direction of propagation of the acoustic wave. Furthermore, we can observe that the damping along the Robin boundary conditions imposed on the sides of the domain depends on the orientation of the nematic director $\mathbf{n}$. In particular, the damping is greater in the region where the nematic director $\mathbf{n}$ is aligned with the direction of propagation of the acoustic wave. This type of phenomena has been predicted from the analysis of the nematic Helmholtz–Korteweg equation [farrell2024b].
The difference between this experiment and the anisotropic propagation of plane waves discussed in the previous subsection is that rather than considering a Gaussian pulse propagating from the centre of the domain, here we consider a planar wave propagating from the top of the domain downwards, and study how the orientation of the nematic director $\mathbf{n}$ affects the speed of propagation of the wave. In particular, while in the previous subsection we considered a nematic director $\mathbf{n}$ with constant orientation throughout the domain, here we consider a nematic director $\mathbf{n}$ with different orientations in different portions of the domain.
Remark 5.1.
The discontinuity in the nematic director field presented in Figure 4 cannot be achieved if the nematic director field is a general solution of the Euler–Lagrange equations associated with the Oseen–Frank energy density, and should be regarded as an approximation of a solution of the Oseen–Frank model exhibiting a sharp transition between two constant orientations.
Figure 4: The anisotropic propagation of a planar acoustic wave in the modified Mullen–Lüthi–Stephen experiment (top) and the corresponding nematic director $\mathbf{n}$ (bottom). The parameters are $k=40$, $\alpha=10^{-4}$ and $\beta=10^{-4}$.
5.4 Tunable Resonators
Resonance occurs when the frequency of an acoustic wave coincides with one of the system’s natural frequencies, corresponding to an eigenfunction of that system. Under these conditions, the wave interferes constructively with itself, leading to a substantial amplification of the wave’s amplitude. This resonant behavior is not only a fundamental concept in wave physics but also a key principle in the design of acoustic resonators. By carefully tuning geometrical and material parameters to match specific eigenfrequencies, engineers and scientists can enhance sound intensity, control wave propagation, and achieve highly efficient energy transfer within acoustic systems.
The nematic Helmholtz–Korteweg equation suggests that a nematic liquid crystal and other nematic materials can be used to design tunable acoustic resonators, since the eigenvalues of the nematic Helmholtz–Korteweg equation depend on the orientation of the nematic director $\mathbf{n}$, as shown in [farrell2024b]. Thus, by changing the orientation of the nematic director $\mathbf{n}$, for example via an external electromagnetic field, it is possible to tune the eigenfrequencies of the system to be closer or further away from the frequency of the incoming acoustic wave, thus controlling the resonant behaviour of the system.
In Figure 5 we present numerical simulations of a tunable acoustic resonator based on the nematic Helmholtz–Korteweg equation. The simulation is obtained solving the following scattering problem for an incoming plane wave $u^{-}$:
$\alpha\Delta^{2}u^{+}+\nabla\cdot\nabla\left(\mathbf{n}^{\top}(\mathcal{H}u^{+})\mathbf{n}\right)-\Delta u^{+}-k^{2}u^{+}$ $=\!0,$ $\text{ in }\mathbb{R}^{2}\backslash\mathcal{R}$, (35)
$u^{+}-u^{-}$ $=0,$ on $\text${ on $\partial\mathcal{R}$
$\alpha\Delta(u^{+}-u^{-})+\beta\mathbf{n}^{\top}\mathcal{H}(u^{+}-u^{-})\mathbf{n}$ $=0,$ on $\text${ on $\partial\mathcal{R}$
$\lvert\partial_{\lvert\mathbf{x}\rvert}u^{+}-iku^{+}\rvert$ $=\mathcal{O}(\lvert\mathbf{x}\rvert^{-\frac{1}{2}}),$ $\lvert\mathbf{x}\rvert\to\infty$,
where $u^{+}$ is the scattered wave and $\mathcal{R}$ is the resonator’s domain. We discretise the associated weak formulation with the Argyris finite element space. We truncate the unbounded domain $\mathbb{R}^{2}\backslash\mathcal{R}$ to a bounded domain via an adiabatic absorbing layer, as discussed in [pml]. In particular, for $\mathbf{n}=(1,0)$ we set $u^{-}=\exp\left(i\mathbf{d}\cdot\mathbf{x}\right)$ with $\mathbf{d}=(0,\sqrt{Re(\lambda_{h}+\epsilon)})$, $\epsilon>0$, where
$\lambda_{h}=22.680155441303704-1.386981204327703\cdot 10^{-7}i$ (36)
is a discrete eigenvalue of the nematic Helmholtz–Korteweg eigenvalue problem (15). Note that we have to add a small perturbation $\epsilon$ to the eigenvalue to ensure that the assumption $k^{2}\not\in\{\lambda_{h}^{(i)}\}$ is satisfied.
((a))
((b))
((c))
Figure 5: (a) The eigenmode associated with the discrete eigenvalue (36) for $\mathbf{n}=(1,0)$. (b) The scattered wave generated by incoming plane wave $\exp\left(i\mathbf{d}\cdot\mathbf{x}\right)$ with $\mathbf{d}=(0,\sqrt{Re(\lambda_{h})})$ for a fixed nematic director $\mathbf{n}=(1,0)$ (c) The scattered wave generated by the same incoming plane for a fixed nematic director $\mathbf{n}=(0,1)$ respectively. Notice the difference in scale: the middle figure has scale $10^{7}$ times larger.
From Figure 5(b) we can observe that if the nematic director $\mathbf{n}$ is aligned with the $x$-axis and the frequency of the incoming plane wave is close to the eigenfrequency of the system for the same nematic director orientation, then the system exhibits resonant behaviour, with a substantial amplification of the wave’s amplitude. It is also possible to observe from Figure 5(c) that if the nematic director $\mathbf{n}$ is aligned with the $y$-axis, then the frequency of the incoming plane wave is no longer close to an eigenfrequency of the system, and the system does not exhibit resonant behaviour. This shows that by changing the orientation of the nematic director $\mathbf{n}$ it is possible to tune the resonant behaviour of the system, thus giving rise to a new class of tunable acoustic resonators.
Acknowledgements
The authors would like to thank Christoph Lehrenfeld and Ilaria Perugia for their valuable comments and suggestions.
Funding
This work was funded by the Engineering and Physical Sciences Research Council [grant numbers EP/R029423/1 and EP/W026163/1], the Donatio Universitatis Carolinae Chair “Mathematical modelling of multicomponent systems”, and the UKRI Digital Research Infrastructure Programme through the Science and Technology Facilities Council’s Computational Science Centre for Research Communities (CoSeC). The second author acknowledges funding through the DFG Grant 432680300 – SFB 1456.
Appendix A Traces
For $\sigma\in H^{1}(\Omega)$ it is well known that the trace $\text{tr}(\sigma):H^{1}(\Omega)\rightarrow H^{1/2}(\partial\Omega)$ is a bounded linear operator and in particular, it is a compact operator from $H^{1}(\Omega)$ to $L^{2}(\partial\Omega)$ [Leoni]. For the normal derivative trace $\text{tr}_{\bm{\nu}}(\sigma)$, the situation is more complicated. In fact, for general $\sigma\in H^{1}(\Omega)$ we have that $\nabla\sigma\in H(\text{curl},\;\Omega)$ for which the normal derivative trace is not well-defined. Yet, the Laplace problem with Neumann boundary condition is well-defined. There, the issue is resolved by the fact that if $\sigma\in H^{1}(\Omega)$ is a solution of the Laplace equation—distributionally in the interior of the domain—then $\nabla\sigma\in H(\text{div},\;\Omega)$ and thus the normal derivative trace $\text{tr}_{\bm{\nu}}(\sigma)$ is well-defined as an element of $H^{-1/2}(\partial\Omega)$ [BBF13, Lem. 2.1.3].
Similarly, for $u\in H^{2}(\Omega)$ the trace and the normal derivative trace are elements of $H^{3/2}(\partial\Omega)$ and $H^{1/2}(\partial\Omega)$, respectively. But, without further information, $\operatorname{tr}\Delta u|_{\partial\Omega}$ is not well-defined for functions $u\in H^{2}(\Omega)$ and we would not be able to make sense of the terms appearing in Section 2. However, we assume that the source term satisfies $f\in L^{2}(\Omega)$, i.e. we assume more regularity that $f\in H^{-2}(\Omega)$. Hence, any solution $u\in H^{2}(\Omega)$ of (1) satisfies $\alpha\Delta^{2}u+\beta\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n}\in L^{2}(\Omega)$ distributionally, implying that $\nabla(\alpha\Delta u+\beta\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n})\in H(\text{div},\;\Omega)$, and thus we would be able to define the normal derivative trace $\operatorname{tr}_{\bm{\nu}}(\alpha\Delta u+\beta\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n})|_{\partial\Omega}$ as an element of $H^{-1/2}(\partial\Omega)$ and a compact trace operator $\operatorname{tr}(\alpha\Delta u+\beta\mathbf{n}^{\top}(\mathcal{H}u)\mathbf{n})|_{\partial\Omega}$ from $H^{2}(\Omega)$ to $L^{2}(\partial\Omega)$. Thus, we can now make sense of (3).
Notice that there is neither a gain in regularity for the solution $u$ of (3), nor any requirements on the regularity of the boundary $\partial\Omega$. In fact, we are only looking for solutions $u\in H^{2}(\Omega)$ of (3) and use the equation in the interior in the domain to define the traces appearing in the boundary conditions. One could separately make elliptic regularity arguments to show that the solution $u\in H^{2}(\Omega)$ actually lives in $H^{4}(\Omega)$, depending on the assumptions that $f\in L^{2}(\Omega)$ and that the domain is sufficiently regular [Grisvard].
Lastly we would like to remark that we could not have looked for a solution $u\in H^{2}(\Omega)$ directly in the subspace $\{v\in H^{2}(\Omega):\alpha\Delta^{2}v+\beta\nabla\cdot\nabla(\mathbf{n}^{\top}(\mathcal{H}v)\mathbf{n})\in L^{2}(\Omega)\}$, since while the space is a closed subspace of $H^{2}(\Omega)$, the $H^{2}(\Omega)$-norm would not be equivalent to the graph norm induced by the operator $\alpha\Delta^{2}+\beta\nabla\cdot\nabla(\mathbf{n}^{\top}(\mathcal{H}\cdot)\mathbf{n})$, which is needed to have completeness of the space.
Appendix B Proof of Theorem 4.7
Proof.
The proof is given in [EG_FE2, Ex. 26.5] and is based on [W86]. We repeat the argument for completeness and make slight adaptations for the setting of T-coercivity.
Let $R_{h}:X\rightarrow X_{h}$ be the discrete solution map associated with $A$, i.e. for all $v\in X$ it holds that $(A(R_{h}v-v),w_{h})_{X}=0$ for all $w_{h}\in X_{h}$. For a bijective operator $T_{h}$, we note that $AT_{h}$ is uniformly coercive if and only if $T_{h}^{-\ast}A$, where $T_{h}^{-\ast}$ is the adjoint of the inverse of $T_{h}$, is uniformly coercive. Thus, there exists a constant $\alpha_{0}>0$ such that for all $v\in X$
$\alpha_{0}\|R_{h}v\|_{X}^{2}$ $\leq|(AR_{h}v,T_{h}^{-1}R_{h}v)_{X}|\leq\|T_{h}^{-1}\|_{\mathcal{L}(X)}\|A\|_{\mathcal{L}(X)}\|v\|_{X}\|R_{h}v\|_{X}.$
Thus, it holds that $\|R_{h}\|_{\mathcal{L}(X)}\leq(\|A\|_{\mathcal{L}(X)}\|T_{h}^{-1}\|_{\mathcal{L}(X)}/\alpha_{0})$ and we conclude that
$\|R_{h}v-v\|_{X}\leq\left(1+\frac{\|A\|_{\mathcal{L}(X)}\|T_{h}^{-1}\|_{\mathcal{L}(X)}}{\alpha_{0}}\right)\inf_{w_{h}\in X_{h}}\|v-w_{h}\|_{X}.$
Therefore, we have that $R_{h}\rightarrow\operatorname{Id}_{X}$ pointwise as $h\rightarrow 0$ which implies that $R_{h}\rightarrow\operatorname{Id}_{X}$ uniformly as $h\rightarrow 0$ on compact sets [EG_FE2, Rem. C.5].
Define $L\coloneqq\operatorname{Id}_{X}+A^{-1}K$ and $L_{h}\coloneqq\operatorname{Id}_{X}+R_{h}A^{-1}K$. Then, we have that $L_{h}\rightarrow L$ in $\mathcal{L}(X)$ due to the uniform convergence of $R_{h}$ to $\operatorname{Id}_{X}$ on compact sets and the compactness of $A^{-1}K$. We note that $L$ is bijective and $L^{-1}L_{h}=\operatorname{Id}_{X}-L^{-1}(L-L_{h})$. Taking $h<h_{1}$ small enough such that $\|L^{-1}(L-L_{h})\|_{\mathcal{L}(X)}\leq 1/2$ ensures that $L_{h}$ is invertible with
$L_{h}^{-1}L=\sum_{k\in\mathbb{N}}\left(L^{-1}(L-L_{h})\right)^{k},$ (37)
such that $\|L_{h}^{-1}\|_{\mathcal{L}(X)}\leq 2\|L^{-1}\|_{\mathcal{L}(X)}\eqqcolon C_{L}$. Consequently, we have that
$\frac{\alpha_{0}}{C_{L}^{2}}\|v_{h}\|^{2}_{X}$ $\leq\alpha_{0}\|L_{h}v_{h}\|^{2}_{X}\leq|(AL_{h}v_{h},T_{h}^{-1}L_{h}v_{h})_{X}|$
$\leq|(A(L-L_{h})v_{h},T_{h}^{-1}L_{h}v_{h})_{X}|+|(\mathcal{A}v_{h},T_{h}^{-1}L_{h}v_{h})_{X}|$
$\leq\|A\|_{\mathcal{L}(X)}\|T_{h}\|_{\mathcal{L}(X)}\|L-L_{h}\|_{\mathcal{L}(X)}\|v_{h}\|_{X}^{2}+|(\mathcal{A}v_{h},T_{h}^{-1}L_{h}v_{h})_{X}|,$
where the second to last step follows from $AL=A(\operatorname{Id}_{X}+A^{-1}K)=\mathcal{A}$. For $h<h_{2}$ small enough, the first term on the right-hand side can be bounded by $\alpha_{0}/2C_{L}^{2}$ such that for $h<h_{0}=:\min(h_{1},h_{2})$ it holds that
$\frac{\alpha_{0}}{2C_{L}^{2}}\|v_{h}\|^{2}_{X}\leq|(\mathcal{A}v_{h},T_{h}^{-1}L_{h}v_{h})_{X}|.$ (38)
Thus, setting $\tilde{T}_{h}\coloneqq L_{h}^{-1}T_{h}$, we conclude that $\mathcal{A}$ is uniformly $\tilde{T}_{h}$-coercive. ∎